A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

$ [483, 490]. $10.00 [385]. $10.50 [344]. $11.95 [363, 505]. $12.00 [409, 488]. $12.95 [532]. $14.95 [348]. $15.50 [343]. $15.90 [387]. $16.95 [495]. $17.00 [392]. $17.95 [358]. $18.50 [521, 529]. $18.95 [522, 510]. $19.00 [511]. $19.95 [513, 514, 525, 539, 536, 524, 492]. $20.00 [486, 489]. $21.00 [462]. $21.25 [457, 457, 477]. $21.50 [519]. $22.50 [326]. $24.00 [346]. $24.50 [347]. $24.95 [389, 461, 364]. $25.00 [327, 485]. $25.95 [487]. $27.50 [359]. $28.00 [407]. $29.25 [481]. $29.50 [324]. $30.50 [460]. $34.50 [325, 476]. $35.50/Dfl [493]. $36.50 [475]. $38 [491]. $41.50 [494, 479]. $45.00 [458]. $52.00 [408]. $55.50 [459]. $58.50 [473].

$65.75 [474]. $73.25 [328]. $80.00 [478]. ; $91.95 [329]. ∫_{0}^{t} f(t) dg(t) [44]. ; 0 [410, 409, 429, 459, 464, 460, 432, 462, 461, ; 431, 521, 493, 513, 514, 487, 522, 525, 537, ; 539, 510, 529, 536, 519, 524, 494, 511, 532]. ; 0-0-444-00238-3 [536]. 0-07-015509-7 ; [464]. 0-12-125550-6 [519]. 0-12-221860-4 ; [521]. 0-12-400550-0 [476]. 0-12-402370-3 ; [529]. 0-12-487950-0 [511]. 0-444-00175-1 ; [510]. 0-444-00179-4 [513]. 0-444-00192- ; [514]. 0-444-00205-7 [522]. 0-444-00210-3 ; [525]. 0-444-00219-7 [532]. 0-444-00221-9 ; [524]. 0-444-00226-X [539]. 0-444-11069-0 ; [432]. 0-444-85111-9 [438]. 0-444-85199-2 ; [494]. 0-444-85208-5 [493]. 0-470-26588-4 ; [537]. 0-471-06118-2 [461]. 0-471-75517-6 ; [429]. 0-471-80344-8 [462]. 0-471-98152-4 ; [482, 431]. 0-7204-0305-7 [478].

0-7204-0368-5 [477]. 0-7204-0387-1 [481]. 0-7204-0406-1 [459]. 0-7204-0437-1 [479]. 0-7204-0474-6 [473]. 0-7204-2806-8 [475]. 0-7204-2829-7 [474]. 0-87815-018-8 [460]. 0-87942-099-5 [487]. 0-901223-69-7 [409]. 0-90285-92-2 [410]. 00 [322, 323].

1 [459, 510, 386, 457]. 1-19 [235]. 105.000 [365]. 10th [488]. 13.00 [327]. 135.000 [361]. 158 [350]. 18th [486]. 190.00 [328]. 1970 [20, 11, 31, 41]. 1971 [61, 51, 70, 79]. 1972 [99, 89, 112, 123]. 1973 [143, 133, 155, 164]. 1974 [181, 172, 190, 200, 474]. 1975 [223, 207, 235, 248, 479]. 1976 [267, 259, 278, 288, 482, 535]. 1977 [333, 309, 299, 317, 345, 435]. 1978 [414, 368, 352, 398, 438, 410]. 1979 [467, 543, 441, 497]. 1C [486]. 1HQ [409]. 1st [483].

2 [410, 465, 463, 461, 494]. 2-604-00033-4 [465]. 2-604-00034-2 [463]. 22.50 [428]. 28.50 [480]. 283 [367].

3 [525, 534, 535, 491, 529, 536, 538, 523]. 3-258-02811-7 [491]. 3-486-21491-8 [523]. 3-486-21551-5 [538]. 3-519-02336-9 [480]. 3-540-08450-9 [535]. 3-540-90302-X [534]. 358 [544]. 3EW [410].

4 [465, 431, 521, 513, 537]. 4C [485, 489]. 4SMC [490].

5 [408, 493, 487, 538]. 54.00 [362].

6 [429, 488, 519]. 60.00 [360]. 65.00 [389]. 6C [484].

7 [407, 409, 464, 522, 491, 532]. 70.- [407]. 75 [430]. 77CH1252 [484]. 77CH1252-6C [484]. 77CH1253 [489]. 77CH1253-4C [489]. 77CH1266 [488]. 77CH1266-6 [488]. 77CH1270 [483]. 77CH1270-8C [483].

77CH1278 [486]. 77CH1278-1C [486]. 77CH1291 [485]. 77CH1291-4C [485]. 77CH1309 [490]. 77CH1309-4SMC [490]. 7th [383].

8 [460, 462, 197, 523]. 8/I [197]. 8C [483]. 8th [329].

9 [535, 524]. 90 [407, 408]. 90-277-0803-7 [407]. 90-277-0852-5 [408]. 92-835-0222-1 [533]. 9th [490].

A/D [364]. Abbildungen [538]. Academic [324, 326, 327, 391, 343, 521, 476, 506, 507, 508, 529, 519, 531, 511]. Accuracy [500, 307]. ACSL [304]. action [170]. Activite [88, 132, 142, 10, 19, 29, 39, 50, 60, 69, 78, 98, 111, 121, 263, 154, 163]. activity [263]. adaption [62]. adaptive [400]. addressable [505]. adjoint [449]. adjustment [125, 421]. Advanced [407, 243]. Advances [511, 396, 435]. after [294]. AGARD [533, 533]. AGARD-CP [533]. Agrawala [343]. Ahcssl [243]. AICA [329]. aide [102, 209, 422]. aided [377]. Aircraft [533, 104, 34]. airfoil [251]. al [425]. alèatoires [418]. algebraic [127, 127]. algébrique [127]. algébriques [127]. ALGOL [538]. ALGOL-Programme [538]. algorithm [301, 176]. algorithmic [393, 395]. algorithms [340, 539]. algorithmus [491]. Allan [387]. along [502]. alternatif [210]. alternating [184]. alternative [210]. altitude [34]. American [432]. Ames [458]. Amplificateur [137]. amplifier [347, 137]. amplifiers [82, 72]. Amsterdam [322, 323, 329, 395, 396, 389, 360, 394, 427, 434, 433, 430, 437, 473, 474, 493, 513, 514, 475, 481, 522, 525, 477, 539, 478, 479, 510, 536, 524, 494, 532, 390]. Analog [30, 44, 211, 422, 187, 45, 442, 103, 3, 126, 169, 14, 156, 113, 4, 5, 262, 115, 83, 254, 53, 92, 106, 252, 127, 73, 64, 480, 280, 492]. analog-digital [254]. Analogies [270]. analogique [103, 3, 126, 156, 113, 280, 148, 115, 92, 422, 106, 127, 73]. analogiques [252]. analogous [285]. Analogrechner [12]. analogue [178, 420, 55, 226, 276, 256, 90, 291, 421]. analysing [420]. Analysis [428, 188, 519, 325, 192, 268, 62, 158, 504, 514, 525, 536, 531, 228, 387]. analytical [369, 415]. analyzer [191]. Ancelle [533]. and/or [255, 313]. Annals [30]. Annotated [456]. any [113]. APAC [13]. Applicability [237, 239]. Application [292, 387, 184, 293, 444, 92, 448, 283, 291, 401, 483, 535, 176, 137, 253, 127, 343]. Applications [415, 310, 3, 14, 445, 483, 485, 495, 530, 363, 459, 485, 514]. applied [158, 204, 428]. approach [384, 388, 460, 202, 194, 443, 232, 393, 532]. approximation [265, 305, 454, 446]. approximations [423, 403, 269, 307, 261]. April [410, 495, 20, 61, 99, 143, 181, 223, 267]. aquifer [184]. architecture [343, 363, 321, 367, 407]. architectures [510]. Aries [534]. arising [282]. arithmetic [2]. arrival [272]. art [167, 521]. arteries [294, 500]. Article [350]. Aspects [321, 367]. assemblies [137]. associant [136]. Association [30, 88, 132, 142, 154]. asynchronized [157]. atmosphere [158, 23]. August [329, 361, 408, 474, 489, 467]. Author [331, 413, 541]. authors [30]. automata [434, 473]. automated [231]. Automatic [348, 125, 421]. Automation [81]. Automatisation [81]. Automatized [193]. aux [293, 137, 209, 283, 114, 127, 530]. Avenue [324, 326]. axial [93].

B [327, 393, 384, 386, 430, 431, 515, 526, 534, 537, 510, 538, 482]. Balaton [479]. Barna [344]. based [305, 135, 513]. basée [135]. Basic [324]. basses [52]. Bauknecht [512]. Bazaraa [518]. be [233]. bed [93]. behalf [409]. behaviour [420]. Beilner [345]. Belgium [361, 365, 474]. Bell [479]. Bellaire [489]. Belmont [385, 387, 426, 436]. Belmot [392]. Benito [491]. Bennett [438]. beobachtung [491]. beobachtungspolitik [491]. Berlin [383, 534, 535, 526, 512]. Bern [491]. Bertsekas [506]. Bessel [403, 256]. between [355]. Bibliographie [171, 180]. bibliographies [456]. bibliographique [7, 16, 26, 36, 47, 57, 66, 75, 85, 95, 108, 118, 129, 139, 151, 160, 198, 206, 222, 234, 247, 257, 277, 286]. Bibliography [189, 266, 257, 277, 297, 314, 349, 366, 397, 411, 540, 439, 496, 286, 308, 330, 466]. Bickford [425]. Biology [535, 473]. Blanc [358]. Blinchikoff [359]. blood [500, 252]. Board [258, 298, 351, 440]. Bogner [380]. Bolyai [328]. Book [324, 322, 325, 326, 323, 327, 328, 329, 367, 395, 396, 358, 361, 377, 378, 390, 407, 346, 347, 391, 348, 362, 389, 343, 359, 392, 379, 344, 380, 393, 408, 345, 360, 363, 381, 382, 383, 394, 410, 365, 384, 385, 386, 387, 409, 364, 388, 428, 435, 456, 425, 429, 459, 465, 464, 427, 460, 434, 426, 432, 433, 462, 463, 438, 430, 458, 461, 436, 437, 457, 431, 473, 492, 521, 533, 474, 483, 484, 485, 486, 493, 505, 513]. Book [514, 475, 476, 487, 506, 507, 508, 515, 481, 522, 525, 526, 534, 477, 537, 539, 509, 517, 518, 535, 478, 479, 488, 491, 495, 510, 527, 528, 529, 530, 536, 480, 489, 519, 524, 531, 538, 490, 494, 511, 512, 523, 482, 532]. Books [384]. Bordeaux [478]. Boston [408]. Boston-USA [407, 408]. bottleneck [273]. boucle [103]. Boulaye [407]. boundaries [14]. boundary [385, 426, 471, 374, 469, 209, 303]. Box [322, 323, 360, 456, 537]. Brameller [387]. Branin [519]. breeze [502]. Bruges [365]. Brunswick [435]. brushless [53]. bubble [211]. bubble-column [211]. building [284]. burning [1]. BWR [444].

C [378, 390, 346, 382, 388, 433, 474, 505, 487, 537, 479, 527, 512]. CA [392, 385, 387, 426, 436]. calcul [101, 113, 114, 149]. calculateur [220, 3, 136, 126, 124, 156, 422, 106, 81]. Calculateurs [210, 209, 145, 224]. Calculation [402, 101, 113, 114, 73]. calculator [220, 3, 136, 126, 124, 156, 422, 106, 81, 73]. calculators [209, 210]. calculatrice [73]. California [456]. Call [316]. can [233]. caractéristiques [73]. Carasso [426]. cardiac [92, 263]. cardiaque [92, 263]. cardiovasculaire [156]. Cardiovascular [294, 156, 168]. care [483, 483]. Carlo [14, 504]. cas [113]. Case [511, 360, 460, 113]. Catastrophe [276]. CCTL [260]. CEA [383]. CEin [491]. Center [358]. certain [12, 124, 214]. certaines [124]. Cesari [507]. CH [491]. chaleur [115]. champ [92, 106]. champs [115, 92, 253]. channels [310]. Chapman [517]. characteristic [73]. Charter [527]. charts [446]. Chebyshev [226]. chemical [382]. Chester [410]. Chicago [485]. Chronique [7, 16, 26, 36, 47, 57, 66, 75, 85, 95, 108, 118, 129, 139, 151, 160, 198, 206, 222, 234, 247, 257, 277, 286]. circuit [325, 347, 264, 465]. circuits [346, 270]. circulation [502, 252]. Clark [379]. class [229]. classes [214]. Cliffs [325, 346, 348, 428]. Cluley [378]. coastline [502]. coefficients [127, 149]. cojxutation [285]. Coles [379]. collocation [353]. Colloquia [328]. Colton [385]. column [211]. combination [185]. combinatorial [509]. combining [136]. Comments [269]. compact [318]. Company [322, 323, 328, 329, 395, 396, 361, 407, 362, 389, 408, 345, 360, 394, 365, 459, 465, 464, 427, 434, 433, 463, 438, 430, 437, 457, 473, 492, 474, 493, 475, 481, 477, 478, 479, 495, 536, 494]. comparative [379]. Comparing [214]. Comparison [230]. complex [192, 337, 502]. complexity [429]. comportant [146, 145]. comprising [146]. computation [348, 428, 254, 445, 30, 425]. computations [399]. Computer [251, 357, 423, 338, 54, 63, 168, 176, 367, 361, 393, 345, 360, 410, 435, 319, 250, 192, 14, 279, 193, 268, 179, 24, 4, 321, 405, 197, 175, 134, 229, 83, 174, 53, 419, 71, 312, 232, 226, 276, 256, 13, 216, 90, 158, 291, 45, 188, 473, 483, 484, 485, 486, 525, 510, 536, 480, 532, 166, 91, 170, 478, 358, 361, 377, 378, 407, 459, 474, 483, 484, 485, 486, 493, 513, 514, 527]. Computer-aided [377]. Computer-based [513]. Computerized [390]. Computers [350, 367, 544, 394, 5, 244, 55, 295, 203, 487, 348]. computing [340, 243, 335, 1, 429]. Computor [32, 82]. Conception [263]. conceptual [460]. concerning [261]. condition [264]. conditions [469, 209, 33, 114, 313]. conductivity [93]. Conference [408, 410, 365, 430, 533, 486, 361, 383, 474, 485, 478, 479, 489, 490]. Congress [329, 396]. conservation [269]. considerations [261]. Constantinides [380]. constituent [337]. Construction [337]. consultations [513]. Content [505]. Continuous [167, 406, 246, 174, 470, 243, 304, 448, 313, 355, 480]. Continuous-system [167, 406, 246, 470]. continuous-time [355]. Contribution [144, 33]. Control [409, 447, 400, 316, 381, 458, 169, 301, 444, 82, 217, 262, 304, 443, 453, 45, 334, 242, 494, 360, 384, 506, 517]. controlled [205, 233]. convenient [254]. convergent [81]. converter [202, 422, 420]. Converters [364]. convertisseur [422]. Corporation [537]. correction [294]. correlated [418]. corrèlès [418]. corresponding [105]. corrupted [144]. Cotton [358]. Councils [456]. counter [401]. counter-current [401]. coupled [231]. CP [533]. Crane [378]. criteria [166]. crossroads [473]. current [401, 336]. cutsets [504]. cybernetic [54]. cylindrical [105].

D [326, 407, 344, 408, 360, 363, 385, 364, 506, 537, 529, 511]. D. [347, 520]. D/A [364]. dans [126, 101, 113, 209, 106, 253]. Dare [71]. Darstellung [80, 2]. d’articles [6, 15, 25, 35, 46, 56, 65, 74, 84, 94, 107, 117, 128, 138, 150, 159]. data [310, 62, 391, 476, 508]. Davis [392, 385, 387, 426, 436, 517]. DDA [175]. December [396, 484, 519, 490, 333, 414, 543]. decision [408, 434, 369]. decisions [544, 443, 501]. decomposition [468]. decoupling [301]. Dekker [329, 382]. delay [42]. delayed [221]. delays [281]. d’éléments [209, 224]. Delft [329]. demagnetizing [402]. Demoyer [384]. Department [435]. dependence [544, 501]. Depeyrot [389, 427]. dépouillement [73]. derivation [52, 269, 52]. descriptive [241]. Design [301, 263, 495, 251, 428, 437, 312, 443, 226, 256, 515, 520, 347, 346, 465, 477]. designed [406]. detection [260]. determinate [101]. determination [253, 296, 101, 253]. determining [14]. deterministic [313, 404]. d’ètude [253]. Development [215, 229, 473]. Developmental [475]. developments [174]. device [52]. Dfl [322, 323, 328, 329, 395, 396, 361, 407, 362, 389, 408, 345, 360, 394, 365, 427, 434, 433, 430, 437, 494]. Dfl/130.00 [408]. diagonal [424]. difference [230, 423, 500, 216, 214, 228, 318]. difference-differential [500]. differences [271, 283, 283]. differential [324, 327, 392, 435, 438, 178, 12, 194, 231, 373, 191, 124, 14, 193, 282, 469, 80, 337, 215, 229, 500, 449, 216, 375, 188, 166, 307, 446]. differential-difference [216]. Differentialgleichungen [80]. Differentialgleichungssysteme [12]. différentialles [124]. diffusion [275, 293, 280, 283, 292, 23, 186]. Digital [169, 197, 380, 462, 191, 238, 280, 444, 285, 254, 71, 295, 312, 448, 183, 528, 480, 346, 487, 515]. digitalen [512]. dimension [102]. dimensional [92]. dimensions [449]. d’impédances [220]. d’intégrales [3]. Direct [469, 22, 320, 336]. direct-current [336]. direction [184]. discharge [185]. discontinuities [373]. discontinuous [33]. Discrete [419, 416, 401, 355, 516, 464]. discrete- [355]. discrete-time [516]. discussion [239]. diskreter [526]. display [179]. dispositif [52]. Distributed [537, 185, 217, 227]. disturbation [384]. division [537]. DM [534, 535, 491, 480, 523]. DM.44 [538]. DMSL [241]. Dobrushin [535]. Documents [8, 17, 27, 37, 48, 58, 67, 76, 86, 96, 109, 119, 130, 140, 152, 161]. Dodes [536]. domains [359]. Dordrecht [407, 408]. Dordrecht-Holland [407, 408]. Dordrecht-Holland/Boston-USA [407]. Dorny [388]. double [450]. doubles [3]. down [187]. Dr [491]. draw [187]. Dreyfus [521]. Drive [392, 385, 387, 426, 436]. drives [202]. driving [33]. dual [443]. d’un [136, 126, 113, 102, 422, 42, 106, 263, 81]. d’une [103, 101]. Durch [12]. Dynamic [104, 336, 472, 544, 499, 268, 179, 262, 416, 501, 521, 225, 103, 391, 476, 506, 508]. dynamical [33, 531, 507]. dynamics [443]. dynamique [103]. Dynamo [82]. Dynamo-electric [82].

EAI [197]. échantilonnage [52]. Economic [362, 481]. economics [406]. Ed [358, 383]. Edited [540]. Editor [327, 328, 329, 396, 361, 394, 365, 435, 430, 493, 478, 494]. Editorial [236, 173, 182, 165, 208, 249, 289, 258, 298, 351, 440]. Editors [392, 345, 434, 426, 438, 535, 519, 511, 479]. Edt [395]. effect [104, 253]. effet [253]. Efficient [373]. effluent [185]. eigenvalues [14]. Einführung [526]. electric [209, 82, 92, 422, 252, 263]. Electrical [409, 178, 342, 135, 176]. electrical-transmission [176]. électrique [209, 92, 422, 263]. électriques [224, 252, 135]. electromagnetic [338]. électromécaniques [270]. Electromechanical [270]. electron [498]. Electronic [342, 126]. électroniques [126]. électrons [498]. element [374, 293, 280, 371, 218, 214, 228, 269]. elements [353, 209, 293, 280, 389, 427, 522]. elevated [319]. elliptic [471, 353, 215, 403, 454, 449]. Elsevier [322, 323, 390, 432]. emphasis [316, 494]. empirical [391, 476, 508]. emploi [224]. Engelwood [428]. Engineering [409, 461, 242, 519, 523, 348]. Engineers [409, 462, 528]. England [410, 409]. Englewood [325, 346, 348]. enthalpy [371]. entries [32]. Entwurf [512]. Environment [533]. equation [178, 12, 373, 113, 280, 102, 292, 311, 127, 113, 280, 102, 127, 326]. equations [327, 435, 438, 194, 231, 353, 124, 14, 282, 469, 80, 337, 148, 215, 229, 500, 449, 354, 283, 216, 375, 127, 188, 176, 166, 307, 446, 318, 149, 124, 293, 127, 324, 392]. equivalence [355]. equivalent [442]. Errata [100]. Erratum [147, 274, 315, 544]. Error [339, 268, 295]. errors [188]. estimation [461, 282, 144, 203, 204, 404, 517, 421]. estuary [232]. états [126]. ETH [491]. Etude [148, 106]. Études [224]. EURO [396]. European [396, 408]. evaluating [442, 399]. evaluation [345, 215, 419, 445, 514, 362, 481, 460]. event [419]. exact [405, 355, 42]. exacte [42]. Examination [341]. experiment [260]. experimentation [377]. exponential [454, 281]. extension [404].

F [324, 395, 347, 533, 491, 519]. F-92200 [533]. factorization [449]. factors [296]. failure [452]. Falkenplatz [491]. Falls [488]. Fan [382]. fast [14, 254, 71]. Faurre [389, 427]. Fearon [392, 385, 387, 426, 436].

Fearon-Pitman [392, 385, 387, 426, 436]. feedback [233, 82]. Feilmeier [394]. ferromagnetic [106]. ferromagnétiques [106]. FFT [402]. fibers [170]. field [92, 55, 106, 338, 402]. fields [115, 92, 253]. Fifth [324, 326]. figures [523]. filter [291, 313, 355, 404]. filtering [380, 359]. filters [448]. Finding [455]. finies [283]. finis [293]. finite [353, 374, 293, 371, 283, 218, 214, 228, 269, 318]. finite-element [293, 269]. Finkel [377]. First [361, 282, 256, 311, 474]. Fisher [360]. fixed [93]. flight [34]. flow [382, 303, 500, 218, 525]. flows [518]. fluides [3]. fluids [3]. following [444]. fonction [101]. fonctionnant [210, 270]. fonctionnement [156]. fonctions [101, 221, 42]. Forest [425]. formulas [191]. formulation [90]. formulations [355]. FORSIM [231]. Fortran [526, 526]. Fortran-Programmpaketes [526]. Foster [505]. foundations [319, 486, 343]. Fourier [339]. fourth [471, 230]. Fractional [265, 403]. France [407, 533]. Franta [524]. Freedman [323]. French [418, 52, 220, 103, 3, 136, 275, 126, 124, 156, 101, 113, 293, 280, 102, 146, 137, 209, 115, 221, 92, 422, 210, 498, 42, 106, 283, 252, 114, 253, 270, 116, 135, 263, 127, 81, 73]. fréquences [52]. frequencies [196]. frequency [359, 52, 228]. Fried [322]. Froebenius [149]. Fuchsiennes [149]. Function [445, 233, 279, 101, 454, 33, 256, 72, 392]. functional [32]. functioning [156]. functions [408, 101, 221, 255, 305, 403, 42, 212, 219, 464]. Fundamental [194]. fuze [62]. Fuzzy [434].

G [361, 407, 343, 380, 360, 365, 434, 432, 473, 474, 475, 509, 479, 511, 523]. G. [350]. gain [72]. Gaithersburg [484, 490]. Garrett [492]. GASP [386]. Gate [409]. Gaussian [534]. Gelenbe [345]. general [90, 539]. generalised [281]. generalized [436]. generated [233]. Générateur [116]. Generation [305, 418, 356, 423, 265, 244, 335, 418]. generations [255]. generator [300, 451, 312, 256, 116, 260, 336]. generators [72]. Georgi [465, 464, 463]. German [12, 80, 2]. Gewald [523]. Gewisser [12]. Ghent [361, 474]. GI [394]. Gilbert [392]. Gilio [480]. Gillet [393]. global [420, 379]. GPSS [526]. GPSS-Fortran [526]. gradient [404]. graphic [179]. gravity [64]. gravity-oriented [64]. Graw [393]. great [294]. Grinold [433]. Grössen [80]. Grosverhor [410]. Groucho [529]. Groun [529]. groundwater [187, 322]. Grove [425]. GU1 [410]. guideways [450]. Guilford [410]. Gupta [434].

H [323, 359, 345, 409, 437, 492, 513, 522, 517, 519]. Haake [523]. Haimes [323]. Haley [430]. Hall [325, 323, 346, 348, 428, 517, 253]. Halstead [522]. Halsted [537]. Hamam [387]. Handbook [364, 347, 527, 325]. Händler [367]. Hardware [300, 310]. harmonic [92]. harmoniques [92]. Haupt [491]. Heady [362, 481]. heat [326, 93, 115]. heat-transfer [93]. heated [338]. heating [357]. Hecht [525]. held [407, 408, 365, 484, 485, 486, 535, 488, 489, 490]. helmsman [372]. hereditary [447]. Herman [475]. Herts [409]. hierarchies [544, 369, 415, 501]. High [410, 416, 72]. High-speed [416]. Higher [175]. Hilfe [526]. Hill [347, 393, 464]. Hnatek [364]. Hoes [358]. Holland [328, 329, 395, 396, 361, 362, 389, 408, 345, 360, 394, 365, 459, 427, 434, 433, 438, 430, 437, 457, 473, 474, 493, 513, 514, 475, 481, 522, 525, 477, 539, 478, 479, 510, 536, 524, 494, 532]. Holland/Boston [407]. Homogeneous [285, 81]. homogènes [81]. Horwitz [384]. Hotel [410]. House [410, 409]. Hubbard [327]. Hungary [479]. Huseyin [519]. Hy [532]. Hybrid [400, 93, 279, 62, 217, 134, 335, 177, 372, 472, 250, 220, 184, 194, 245, 136, 124, 193, 237, 268, 282, 238, 80, 101, 179, 24, 405, 197, 264, 83, 254, 244, 2, 243, 304, 232, 13, 216, 203, 114, 158, 23, 81, 188, 480, 91, 275, 170, 193, 301]. hybrid/digital [238]. Hybride [80, 220, 136, 275, 124, 101, 114, 81]. hybriden [2]. hydrodynamic [213]. hydrologic [183]. hydrothermal [371]. hyperbolic [230, 214, 176, 318].

I.S.A [381]. IA [458]. Ibragimov [534].

IBSN [438]. Identification

[511, 457, 250, 21, 291, 302, 313, 183]. IEE [409]. IEEE [358, 483, 484, 485, 486, 487, 488, 489, 490]. IFIP [361, 383, 365, 474, 479]. IFORS [430]. II [396, 310, 71, 219, 383, 435]. III [327]. IMACS [394, 435, 494, 316]. IMACS-Symposium [316]. Image [350, 306]. impedance [220, 125]. Implementation [186, 194, 240, 515]. implementations [195]. Implicit [227, 184, 271, 318]. implicitly [231]. Importance [452]. imposition [114]. Improperly [426]. improved [4]. incremental [175, 55, 445]. indefinis [92]. indefinite [92]. Index [40, 122, 199, 287, 331, 332, 413, 412, 541, 542]. Indiana [537]. indicates [34]. indirekter [491]. inductance [43]. induction [205]. inductively [338]. industrial [262, 443]. infants [294]. inflight [62]. Information [408, 460, 408]. Informations [9, 18, 28, 38, 49, 59, 68, 77, 87, 97, 110, 120, 131, 141, 153, 162]. inherent [255]. Initial [264, 405, 33, 313, 228]. insertion [264]. Institute [407]. Institution [409]. Instituts [491]. instrument [34]. integer [256]. integral [385, 124, 311]. intégrales [124]. integrals [3, 403, 454]. integrated [72]. integration [436, 373, 455]. integrators [175]. integro [124, 124]. integro-differential [124]. intégro-différentialles [124]. interacting [535]. Interaction [246, 535]. interactive [192]. Interconnected [409]. interdisciplinary [201]. interface [187, 188]. Interfacing [377, 378]. internal [115]. International [30, 425, 464, 430, 478, 489]. internes [115]. Interpolators [177]. Introduction [209, 516, 539, 463, 344, 380, 393, 537]. inventory [443]. inversion [499, 279]. inverter [417]. Iowa [458]. IPC [410]. ISBN [407, 408, 410, 409, 429, 459, 465, 464, 460, 432, 462, 463, 461, 431, 473, 521, 533, 474, 493, 513, 514, 475, 476, 487, 481, 522, 525, 534, 477, 537, 539, 535, 478, 479, 491, 510, 529, 536, 480, 519, 524, 538, 494, 511, 523, 482, 532]. isothermal [211]. itératif [114, 81]. Iterative [471, 503, 405, 229, 81, 114, 477].

J [322, 377, 378, 348, 359, 379, 408, 381, 383, 425, 429, 431, 515, 509, 518, 479, 527, 528, 530, 512, 532]. J. [516]. Janos [328]. January [11, 51, 89, 133, 172, 207, 259]. Jarvis [518]. Jensen [325]. Jersey [435]. John [537, 482]. Jolla [456]. July [361, 430, 474, 31, 70, 112, 155, 190, 235, 278]. June [456, 309, 368].

Kalman [291, 313, 355, 404]. Kalman-type [313]. Kashyap [391, 476, 508]. Kaufman [347]. Kaufmann [529]. kidney [401]. Kilmann [437]. kind [256, 311]. kit [62]. Kohlaas [512]. Kozesnik [408]. kraft [381]. Kryukov [535]. Kussmaul [436].

L [325, 329, 391, 382, 384, 385, 544, 514, 476, 507, 535]. Lafayette [537]. Lainiotis [511]. Lake [479]. land [502]. Lane [358]. language [406, 240, 241, 254, 243, 304, 183]. Languages [238, 325, 246, 167, 473, 475]. Laplace [311]. Large [503, 433, 334, 531]. l’Association [10, 19, 29, 39, 50, 60, 69, 78, 98, 111, 121, 163]. lattice [212, 219]. Law [521]. layer [303]. layered [218]. leak [260]. Lee [346]. lengths [340]. Lewin [407]. Lexington [384, 384]. librations [64]. limites [209, 114]. limits [114]. Lindenmayer [473]. line [378, 301, 71, 21, 302]. linéaire [275]. linéaires [293, 115, 81]. Linear [32, 503, 324, 43, 12, 282, 293, 444, 255, 33, 45, 320, 81, 212, 219, 261, 509, 517, 518]. Lineare [12]. linéaries [283]. lines [176]. liquid [336]. liquid-metal [336]. Liu [515]. load [444]. load-following [444]. loads [342]. Locally [535]. locked [63]. logic [346, 459, 520]. logical [453]. London [476, 482, 391, 379, 380, 517, 520]. long [433]. loop [400, 103]. loops [83, 63]. Lösender [12]. low [196, 52]. lower [23]. lowest [14]. LSI [406]. Ltd [409].

M [395, 396, 347, 348, 362, 389, 394, 387, 427, 434, 521, 481, 522, 525, 539, 517, 518, 478, 495, 528]. M.I.S. [350, 306]. Mac [393]. machine [157, 53]. machinery [319]. Machtey [539]. macro [241]. maglev [450]. magnetic [106]. magnétique [106]. magnetohydrodynamic [336]. man [383, 192, 478]. man-model [192]. manpower [433, 433]. MAPS [437]. Marcel [382]. March [394, 409, 492, 535, 299, 352, 441]. Marios [478]. marketing [169]. Markov [535, 300, 310, 491]. Marshall [433]. martingales [436]. mass [401]. Masson [530]. Mathematica [328]. Mathematical [457, 319, 501, 290, 399, 544, 432, 462, 528, 529]. Mathematics [350, 367, 544, 370, 473, 394]. Matric [32, 82]. Max [530]. McGlynn [363]. McGraw [347, 464]. McGraw-Hill [347, 464]. McNamee [325]. MD [484, 490]. Mean [307, 340]. means [233, 126, 106, 313, 402]. measurement [514, 125]. measurer [196]. mécanique [3, 156]. mechanical [105, 156]. mechanics [3]. medical [483, 513]. meeting [408]. Mehra [511]. message [144]. mesures [530]. metabolic [337]. metal [336]. Meteosat [306]. METEOSTAT [350]. Method [253, 385, 374, 424, 282, 293, 4, 405, 22, 221, 195, 420, 55, 227, 127, 218, 284, 1, 318, 468]. Mèthode [253, 293, 115, 221, 127]. Méthodes [283, 530]. Methods [442, 503, 395, 392, 435, 438, 471, 341, 469, 215, 229, 354, 376, 283, 291, 214, 296, 514, 166, 228, 399, 432]. Meyer [279]. MI [489]. Michel [531]. Micro [488, 527]. Micro-computer [527]. microcomputers [344]. microprocessor [292]. microprocessors [492, 344, 363, 495]. Microprogrammable [510]. microprogramming [343, 488, 488]. Miller [531]. minicomputer [302, 479]. minicomputers [240, 479, 377, 492]. Minimization [212, 219]. minimum [185]. missile [158, 62]. mobile [101]. Modal [195, 443, 186]. model [273, 192, 169, 170, 193, 371, 451, 304, 453, 284, 372, 125, 421]. model-building [284]. Modèles [135]. Modeling [383, 457, 369, 334, 201, 362, 384, 462, 544, 319, 303, 290, 502, 501, 528, 428, 381]. Modelle [538]. Modelling [433, 316, 431, 419, 474, 481, 494, 482, 345]. modells [491]. Models [529, 391, 379, 388, 460, 337, 448, 135, 401, 183, 476, 508, 433, 382]. modern [244, 463]. modifications [195]. modular [191]. Montages [137]. Monte [14, 504]. Moscow [535]. motors [205]. mouvement [73]. movement [73]. moyen [126, 106]. Multiobjective [323]. multiple [459]. multiple-values [459]. multiplexing [264]. multiplier [233]. multistage [443]. Multivariable [360]. München [538, 523]. Munich [394]. Murty [509]. MYCIN [513].

N [387, 388, 434, 531]. N.J [358, 428]. N.Y [521, 513, 514, 522, 525, 537, 539, 510, 529, 536, 519, 524, 511, 532]. NATO [407, 533]. need [187]. negative [146, 137]. negatives [145, 146, 137]. Netherlands [322, 323, 329, 360, 459, 438, 457]. Network [146, 463, 220, 136, 126, 272, 102, 209, 210, 518]. networking [484, 358, 484]. networks [280, 292, 114, 504, 493]. Neuilly [533]. Neuts [395]. New-York [482]. Niagara [488]. Nicholson [409]. NJ [325, 346, 348]. NL [345]. No [533]. noise [310, 144, 291]. nombres [418]. non [43, 275, 282, 293, 444, 115, 255, 283, 45, 532, 1]. non-linear [43, 282, 444, 255, 45]. non-ohmic [255]. non-steady [1]. non-technical [532]. nonlinear [275, 193, 293, 115, 283, 372]. Normalized [449]. normative [437]. North [328, 329, 395, 396, 361, 362, 389, 345, 360, 394, 365, 459, 427, 434, 433, 438, 430, 437, 457, 473, 474, 493, 513, 514, 475, 481, 522, 525, 477, 539, 478, 479, 510, 536, 524, 494, 532]. North-Holland [328, 459, 457, 473, 474, 493, 513, 475, 481, 477, 478, 479, 510, 494]. Nostrand [505]. note [376, 417]. novel [374]. November [396, 485, 486, 497]. nuclear [444]. Number [456, 312]. numbers [418]. Numerical [213, 424, 502, 311, 327, 271, 136, 253, 455, 307, 446, 283, 327, 438, 536]. numérique [136, 280, 253]. numériques [283]. NY [488].

October [483, 486, 488, 41, 79, 123, 164, 200, 248, 288, 317]. ohmic [255]. Oldenburg [538, 523]. On-line [301, 302, 378, 71, 21]. one [102]. open [400]. operating [210, 270]. operation [390, 378]. Operational [430, 347, 137, 430]. opérationnel [137]. Operations [491, 328, 396, 393]. operators [385]. Optimal [340, 185, 301, 444, 45, 455, 417, 12]. Optimale [12]. optimierung [491, 12].

Optimization [457, 191, 262, 323, 383, 12, 24, 22, 63, 296, 388]. order [271, 282, 230, 405, 175, 256]. ordinary [373, 375]. organizations [433]. oriented [393, 245, 2, 64]. Orthogonal [353]. other [310]. output [341]. Oxford [394, 520].

P [325, 358, 389, 425, 427, 426, 431, 492, 506, 539, 482]. P. [540]. P.O [322, 323, 360, 456, 537]. package [229]. Pages [20, 11, 31, 41, 61, 51, 70, 79, 99, 89, 112, 123, 143, 133, 155, 164, 181, 172, 190, 200, 223, 207, 235, 248, 267, 259, 278, 288, 333, 309, 299, 317, 414, 368, 352, 398, 438, 467, 543, 441, 497, 324, 322, 325, 326, 323, 327, 328, 329, 395, 396, 361, 377, 378, 390, 346, 391, 362, 389, 359, 392, 379, 380, 393, 345, 360, 363, 381, 394, 384, 385, 386, 387, 364, 388, 428, 435, 429, 459, 465, 464, 427, 460, 434, 426, 432, 433]. pages [462, 463, 430, 458, 461, 436, 437, 457, 431, 473, 492, 521, 474, 483, 484, 485, 486, 493, 505, 513, 514, 475, 476, 487, 506, 507, 508, 515, 481, 522, 525, 526, 534, 477, 537, 539, 517, 518, 535, 478, 479, 488, 491, 495, 510, 527, 528, 529, 530, 536, 480, 489, 519, 524, 531, 490, 494, 511, 512, 523, 482, 532, 516, 520, 382, 383, 456, 425]. Panel [239]. papers [316, 533]. parabolic [400, 216, 468]. Parallel [354, 448, 320, 394, 505, 489, 394, 489]. parallelism [367, 321]. Parameter [204, 183, 217, 13, 203, 63, 227, 404, 461]. parameters [250, 194, 313, 421]. parametric [127, 81]. paramétriques [127, 81]. Paris [530]. Part [383, 300, 310, 212, 219]. partial [324, 327, 435, 178, 194, 231, 14, 282, 469, 215, 229, 449, 216, 188, 166, 307, 446]. particle [101]. particule [101]. partition [101, 455]. PASCAL [470]. pathological [209]. pathologiques [209, 224]. Pattern [91]. PDE [503, 468]. PDE-Related [503]. PDEL [183]. PDP [197]. Pegden [537].

Peled [515]. pendulum [105]. Peregrinus [409]. Performance [5, 44, 419, 514]. performances [345]. period [196]. periodic [469]. perspective [480]. Peter [409]. Petrocelli [527]. Petrocelli/Charter [527]. Pfadler [523]. phase [205, 105, 356, 451, 63]. phase-controlled [205]. phase-locked [63]. phenomena [275, 498]. phenomènes [275, 498]. Physical [166, 253, 125]. physiological [168, 204]. physiques [253, 530]. piecewise [212, 219, 261]. Piloted [533]. Piscataway [358]. Pitman [392, 385, 387, 426, 436]. Pittsburg [381]. PL [386]. PL/1 [386]. Planar [64]. planning [433]. plants [444]. plaques [253]. plates [253]. PLL [63]. plotting [55]. point [374]. policy [362, 460, 443, 481]. pollution [322]. polynomial [305, 42, 226]. polynomiales [42]. polynomials [212, 219]. Pontryagin [185]. Porat [344]. posed [426]. potential [170, 101, 113]. potentiel [101, 113]. potentiometer [197]. pour [52, 42, 114, 135]. power [390, 425, 342, 444, 265, 403, 453, 458]. pp [348, 343, 344, 509]. practical [387, 452, 261]. Prague [408]. Prandtl [279]. Precision [271, 103, 103]. Prekopa [328]. Prentice [325, 346, 348, 428]. Prentice-Hall [428]. presence [115, 291]. Press [324, 326, 327, 391, 343, 410, 460, 458, 521, 476, 487, 506, 507, 508, 537, 529, 519, 531, 511, 520]. pressure [371]. pressure-enthalpy [371]. Price [493]. Principes [136]. principle [185]. principles [269, 136, 480]. priorities [544, 501]. Pritsker [386, 537]. probabilistic [416]. probability [395]. problem [471, 245, 405, 217, 33, 2, 214, 417, 519]. problem-oriented [2]. problèmes [224]. problemorientierte [2]. Problems [520, 385, 213, 192, 374, 230, 369, 218, 228, 225, 426]. Procal [242]. procédés [252].

Procedure [149, 184]. Procédures [146, 146, 449]. Proceedings [327, 329, 396, 361, 394, 365, 456, 430, 533, 483, 484, 485, 535, 478, 479, 488, 489, 490, 494, 474, 383, 410]. process [21, 114, 81, 524]. Processes [408, 300, 310, 262, 2, 252, 535, 186, 408, 434, 534]. processing [489, 516, 515]. processors [442, 487]. processors. [505]. Processus [114, 81]. production [381, 443]. production-inventory [443]. Program [244]. programmable [285, 116, 116]. Programme [538, 21]. programmed [13]. Programming [457, 470, 432, 193, 90, 506, 518, 521, 509]. Programmpaketes [526]. programs [525]. Progress [328]. propagation [339]. propellants [1]. proposal [13]. propres [224]. Providence [486]. Publ [491]. publication [30]. Publications [483, 484, 485, 486, 488, 489, 490]. Published [407, 408, 410, 409, 521, 493, 513, 514, 522, 525, 534, 539, 535, 510, 529, 536, 519, 524, 538, 511, 523, 532]. Publishers [392, 385, 387, 426, 436]. Publishing [322, 323, 328, 329, 395, 396, 361, 407, 362, 389, 408, 345, 360, 394, 365, 459, 465, 464, 427, 434, 433, 463, 438, 430, 437, 457, 473, 492, 474, 493, 475, 481, 477, 537, 478, 479, 495, 536, 494]. puits [101, 113]. pulses [310]. Purkinje [170]. Pushchino [535].

quadratic [296]. quadrature [191]. Qualitative [531]. quality [232]. quelconque [113]. queueing [272, 134].

R [325, 358, 391, 362, 392, 380, 363, 384, 386, 387, 364, 428, 435, 433, 438, 437, 476, 508, 481, 477, 535, 479, 529, 524, 531, 538, 511, 523, 532]. Ram [350]. random [408, 418, 312, 534]. range [433]. Rao [391, 476, 508]. Raphael [407]. rates [272]. Rauscher [343]. reactor [211]. reactors [382]. real [302]. Réalisation [135, 263]. realization [263, 135]. Realzeitrechner [538]. Realzeitrechner-Modelle [538]. Rechenautomaten [512]. Rechenprozessen [2]. recognition [91]. Recommendations [30]. reconstructors [177]. record [486]. rectangulaires [253]. rectangular [253]. recus [86, 119, 8, 17, 27, 37, 48, 58, 67, 76, 96, 109, 130, 140, 152, 161]. reduction [62, 295]. redundancy [316, 494]. reentering [158]. reflection [498]. réflexion [498]. regime [275, 210, 270, 275, 210, 270]. Reidel [407, 408]. Reinhold [505]. Related [503, 300, 310, 168]. reliability [529]. reliable [337]. Remoortere [540]. remote [453]. repeated [337]. répétitif [113]. repetitive [14, 113]. reports [34]. representation [105, 80, 92, 2, 92]. reprints [487]. Research [491, 396, 393, 430, 201, 328, 396]. réseau [220, 136, 126, 102, 209]. reseaux [145, 280, 146, 210, 224, 114]. reservoirs [371]. Resistance [210]. resistances [136, 102, 146, 137, 145, 136, 126, 102, 146, 137, 210]. resistivities [255]. resistors [126]. Résolution [124, 102, 103, 113, 81, 103, 113, 81, 124, 102]. resource [361]. Resources [460, 323, 361, 474, 365]. Reston [492, 495]. results [301, 281, 399]. Resumes [15, 25, 35, 46, 56, 65, 74, 84, 94, 107, 117, 128, 138, 150, 159]. Resums [6]. retardateur [42]. retardées [221]. reversible [422, 422]. Review [324, 322, 325, 326, 323, 327, 328, 329, 350, 367, 395, 396, 358, 361, 377, 378, 390, 407, 346, 347, 391, 348, 362, 389, 343, 359, 392, 379, 344, 380, 393, 408, 345, 360, 363, 381, 382, 383, 394, 410, 365, 384, 385, 386, 387, 409, 364, 388, 428, 435, 456, 425, 429, 459, 465, 464, 427, 460, 434, 426, 432, 433, 462, 463, 438, 430, 458, 461, 436, 437, 457, 431, 473, 492, 521, 533, 474, 483, 484, 485, 486, 493, 505]. Review [513, 514, 475, 476, 487, 506, 507, 508, 515, 481, 522, 525, 526, 534, 477, 537, 539, 509, 517, 518, 535, 478, 479, 488, 491, 495, 510, 527, 528, 529, 530, 536, 480, 489, 519, 524, 531, 538, 490, 494, 511, 512, 523, 482, 532, 446]. RI [486]. riding [450]. river [273]. Roubens [396]. Rozanov [534]. Rozenberg [473, 475]. rue [533]. rule [455]. Russak [378]. Russian [534]. Rutgers [435].

S [390, 346, 379, 521, 493, 525, 518]. S. [516]. S.C.S [456]. Saaty [544]. Salazar [487]. Salisbury [510]. sampling [52, 452]. sang [252]. Saphorin [465, 464, 463]. Saridis [434]. satellites [64]. Savage [429]. Savulescu [390]. scale [531]. Scaling [225, 12, 4, 90]. Schad [362, 481]. schemes [230]. Schmidt [526]. Schoemaker [493]. Scholarity [425]. School [535]. School-Seminar [535]. Schriften [491]. Schrödinger [113, 102]. Science [410, 435, 519, 459, 486, 480, 461, 473, 486, 522, 536]. sciences [370, 532]. Scientific [322, 323, 348]. scientists [462, 528]. scope [444]. sea [502]. Seborg [360]. second [396, 271, 405, 478]. Seepage [218]. Seine [533]. seiten [538]. selected [487]. selection [194, 215]. self [43, 449]. self-adjoint [449]. self-inductance [43]. semi [300, 310]. semi-Markov [300, 310]. semibatch [211]. semiconductor [338]. semidiscretization [376]. Seminar [535]. Separate [255]. September [407, 365, 478, 479, 494, 398]. sequence [116, 116]. serial [194]. Series [409, 456, 348, 149]. Service [358, 383, 242]. Services [425]. ses [3]. sets [231]. setting [13]. Seventh [408, 430]. SG1 [409]. Shah [348]. sheets [106]. Shortliffe [513]. signal [144, 487, 515, 530, 516]. signals [52]. signaux [52]. simple [500]. simulated [43]. simulateur [263]. simulating [371]. Simulation [350, 367, 456, 544, 196, 205, 275, 126, 105, 185, 156, 272, 280, 240, 157, 34, 53, 422, 450, 498, 42, 252, 260, 533, 526, 537, 512, 472, 400, 316, 361, 348, 379, 410, 460, 462, 319, 406, 340, 202, 300, 310, 184, 294, 273, 271, 93, 246, 193, 237, 301, 341, 424, 268, 357, 238, 179, 342, 211, 241, 452, 217, 104, 262, 115, 290, 221, 195, 174, 254, 470, 419, 71, 416, 243, 420, 304, 167, 338, 54, 63, 158, 168, 227, 336, 375, 23, 183, 239, 64, 474, 528, 480, 490]. simulation [1, 176, 201, 372, 225, 310, 126, 498, 456, 438, 431, 493, 524, 482, 329, 386, 490, 494]. simulation-language [254]. simulations [373, 512]. Simulations-technik [512]. simulator [444, 263, 350, 306]. Simultaneous [296, 255]. sinewaves [356, 451]. single [450]. sinusoidal [270]. Sippl [527]. situation [273]. size [80]. Skalierung [12]. SLAM [537]. slurry [211]. small [348, 71]. Smith [528]. social [370, 532, 437]. societal [192]. Societatis [328]. software [343, 254, 485, 522, 479, 523]. soils [218]. solar [357]. solid [1]. solides [126]. solids [126]. Solution [127, 327, 471, 213, 194, 231, 282, 469, 337, 134, 449, 232, 292, 216, 311, 188, 385]. solutions [520, 149]. solve [318]. solved [218]. solvers [320]. Solving [503, 375, 468, 405, 229, 354]. Some [281, 261, 230, 262]. sommaire [40, 287, 199]. son [92]. Sons [537, 482]. sources [115]. South [409]. space [388, 449]. span [450]. Sparse [503]. Sparsity [387]. Spatially [401, 185]. special [316, 494]. spectral [341]. spectrum [290]. speech [144]. speed [442, 416]. Spillers [477]. splitting [424]. Springer [383, 526, 534, 535, 512]. Springer-Verlag [383, 526, 534, 535, 512]. squared [307]. St [407, 465, 464, 463]. St-Saphorin [464, 463]. stabilisation [146]. Stabilite [145, 103, 137]. Stability [472, 137, 446, 103, 458]. stabilization [146]. stabilize [72]. Stabilized [356, 233, 451]. State [458, 167, 1]. state-of-the-art [167]. states [126]. static [202, 422, 420]. stationary [115]. stationnaires [115]. statique [422]. statistical [408]. Statisticians [408]. steady [1]. steam [260]. steuerungs [491]. Stevenage [409]. stiff [424]. stimulation [365]. stochastic [391, 295, 312, 504, 476, 506, 508, 517, 436, 457]. Stockholm [396]. Stone [426]. Stout [347]. Street [410]. structural [477]. Structure [409]. structures [175, 285]. Studienbücher [480]. Studies [457, 511]. Study [407, 360, 294, 24, 226, 256, 187, 168, 253, 23, 313, 529, 106, 379, 460]. Subject [332, 412, 542]. supporting [319]. surface [105]. surgical [294]. Surrey [410]. survey [167, 320]. Svobodova [514]. switching [464]. Switzerland [465, 464, 463]. Sydney [482]. Symposium [435, 327, 394, 483, 484, 486, 494, 316]. synchronisation [245]. synchronous [157, 53]. synspade [327]. synthesis [448]. System [421, 389, 384, 425, 427, 458, 406, 245, 192, 231, 246, 156, 357, 342, 217, 174, 470, 244, 243, 443, 167, 453, 203, 291, 320, 447, 302, 313, 480, 225, 365, 511]. system-parameter [203]. système [156]. Systemen [512]. Systèmes [135, 81]. Systemic [252]. systémique [252]. Systems [503, 158, 526, 537, 472, 400, 316, 361, 382, 387, 437, 457, 12, 499, 193, 424, 268, 179, 452, 264, 290, 134, 33, 419, 416, 448, 335, 54, 168, 204, 135, 81, 183, 334, 492, 474, 475, 535, 494, 409, 323, 329, 361, 390, 345, 381, 507, 529, 531].

T [323, 362, 343, 382, 428, 433, 544, 475, 481]. tôles [106]. Tabellen [538]. table [187]. tables [523]. teaching [284]. technical [532]. technik [512]. technique [14, 216, 404, 186]. techniques [251, 383, 295, 530, 533, 325]. Technology [363, 410, 437]. temperature [338]. test [301]. testing [5]. Teubner [480]. texts [30]. their [179, 337, 448, 535]. theorem [135]. Théorème [135]. theoretic [392]. Theory [459, 312, 408, 428, 465, 463, 437, 443, 276, 369, 521, 515, 539, 389, 427, 431, 482]. thermal [115, 135]. thermiques [115, 135]. thermodynamic [101]. thermodynamiques [101]. Thévenin [135]. third [335]. Thompson [481]. Thomson [362]. Thrall [362, 481]. three [356, 451, 449, 92]. three-dimensional [92]. three-space [449]. time [359, 544, 272, 501, 302, 313, 355, 516]. Tokyo [430]. tools [239]. Toom [535]. topology [453]. Toronto [482]. traitement [530]. Transaction [408]. Transactions [407]. Transfer [233, 93]. transforms [339]. transient [93]. transients [425]. Translated [534]. transmission [498, 176]. transport [401]. transposition [294]. trapezoidal [455]. treatment [185]. très [52]. Tretter [516]. Treves [324]. tridimensionnels [92]. Troch [494]. true [34]. turbulence [104]. Turbulent [303, 23]. two [374, 305]. two-point [374]. two-variable [305]. type [175, 313, 404].

U [436]. U.S.A [358, 435]. UKSC [410]. Ultra [254]. Ultra-fast [254]. unconfined [184]. unit [44, 226]. University [458, 435, 520]. unstructured [369]. US$12.00 [484]. US$25 [537]. US$38.95 [345]. USA [324, 326, 407, 408, 345, 428, 456, 425, 458, 492, 537, 495, 392, 385, 387, 426, 436]. Use [503, 291, 454, 203, 504, 101, 113]. User [364, 245]. user-oriented [245]. uses [370]. Using [304, 348, 170, 14, 185, 341, 102, 209, 422, 420, 270, 302, 183, 64, 455, 91, 532]. utilisant [270]. Utilisation [101, 113].

V [326, 535]. valeurs [224, 73]. validity [399]. value [385, 426, 471, 374, 405, 228]. values [459, 73]. Vansteenkiste [361, 365, 474]. variable [280, 305]. variables [280]. various [340, 14]. varying [272]. vector [388]. vehicle [73]. vehicles [450]. véhicules [73]. Verlag [383, 526, 534, 535, 538, 512]. Veronis [495]. versatile [300]. very [52]. vibrating [319]. Vichnevetsky [435, 438]. Vienna [494]. view [524]. Virginia [492, 495]. vol [457]. Volume [408, 456, 331, 332, 413, 412, 541, 542]. volumes [328]. vortex [303, 104].

W [325, 323, 367, 358, 407, 438, 477, 524, 523]. wake [104]. Walke [538]. warhead [62]. Washington [460, 483]. waste [185]. water [323, 361, 362, 365, 384, 232, 474, 481]. weather [271]. well [101, 113]. Wen [382]. West [537]. Widder [326]. Wiley [377, 359, 379, 344, 380, 363, 386, 364, 388, 429, 462, 461, 431, 515, 537, 509, 518, 528, 482, 516]. Williams [381]. Wilnatch [392]. winter [490]. within [444]. without [337]. Working [361, 365, 474]. workshop [488]. world [193].

X [534, 539]. XIX [367, 331, 332, 350]. XX [413, 412]. XXI [541, 542, 544].

Yeh [428]. York [324, 326, 329, 347, 389, 343, 359, 344, 363, 382, 394, 364, 429, 432, 462, 461, 431, 473, 493, 505, 475, 476, 481, 477, 509, 478, 494, 482, 395, 396, 377, 378, 393, 386, 388, 427, 434, 433, 430, 437, 506, 507, 508, 515, 518, 527, 528, 531, 516]. Young [386, 539].

Zehnder [512]. Zeigler [431]. Ziegler [482]. Zissos [520]. Zoutendijk [432]. Zuerev [359]. zur [491]. Zürich [491].

References

Vichnevetsky:1970:CMS

[1] R. Vichnevetsky. A computing method for the simulation of non-steady state burning of solid propellants. Mathematics and Computers in Simulation, 12(1):2–8, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0378475470800192.

Mannshardt:1970:PDH

[2] Reinhold Mannshardt. Eine problemorientierte Darstellung von hybriden Rechenprozessen. (German) [A problem-oriented representation of hybrid arithmetic processes]. Mathematics and Computers in Simulation, 12(1):9–20, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0378475470800209.

Brun-Gabriel:1970:CAI

[3] J. M. Brun-Gabriel. Un calculateur analogique d’intégrales doubles et ses applications à la mécanique des fluides. (French) [An analog calculator for doubles integrals and its applications in the mechanics of fluids]. Mathematics and Computers in Simulation, 12(1):21–26, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0378475470800210.

Hall:1970:IMA

[4] C. R. Hall and S. J. Kahne. An improved method for analog computer scaling.

Mathematics and Computers in Simulation, 12(1):27–32, January 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800222.

Holst:1970:PTA

[5] Per A. Holst. Performance testing of analog computers. Mathematics and Computers in Simulation, 12 (1):33–38, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800234.

Anonymous:1970:RAa

[6] Anonymous. Resums d’articles. Mathematics and Computers in Simulation, 12(1):39–40, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800246.

Anonymous:1970:CBa

[7] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 12(1):41–50, January 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800258.

Anonymous:1970:DRa

[8] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 12(1):51–56, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547080026X.

Anonymous:1970:Ia

[9] Anonymous. Informations. Mathematics and Computers in Simulation, 12(1):57–61, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800271.

Anonymous:1970:ALa

[10] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 12(1):61–62, January 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800283.

Anonymous:1970:PJa

[11] Anonymous. Pages 2–62 (January 1970). Mathematics and Computers in Simulation, 12(1):??, January 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Behringer:1970:OSG

[12] Fred Alois Behringer. Optimale Skalierung Gewisser auf dem Analogrechner zu Lösender Differentialgleichungssysteme Durch Lineare Optimierung. (German) [Optimal scaling of certain differential equation systems by linear optimization]. Mathematics and Computers in Simulation, 12(2):64–74, April 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800015.

Paker:1970:PHC

[13] Y. Paker. A proposal for a hybrid computer with programmed parameter setting — APAC. Mathematics and Computers in Simulation, 12(2):75–80, April 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800027.

DAquanni:1970:MCT

[14] Richard T. D’Aquanni. A Monte Carlo technique using a fast repetitive analog computer for determining lowest eigenvalues of partial differential equations for various boundaries, with applications. Mathematics and Computers in Simulation, 12(2):81–90, April 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800039.

Anonymous:1970:RAb

[15] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 12(2):91–92, April 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800040.

Anonymous:1970:CBb

[16] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 12(2):93–100, April 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800052.

Anonymous:1970:DRb

[17] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 12(2):101–106, April 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800064.

Anonymous:1970:Ib

[18] Anonymous. Informations. Mathematics and Computers in Simulation, 12(2):107–113, April 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800076.

Anonymous:1970:ALb

[19] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 12(2):114–115, April 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800088.

Anonymous:1970:PA

[20] Anonymous. Pages 64–115 (April 1970). Mathematics and Computers in Simulation, 12(2):??, April 1970. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Madden:1970:PLP

[21] Barry G. Madden and Dennis A. Mason. A programme for on-line process identification. Mathematics and Computers in Simulation, 12 (3):118–121, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847547080009X.

Humo:1970:DMO

[22] Emir A. Humo. On a direct method of optimization. Mathematics and Computers in Simulation, 12 (3):122–129, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800106.

Shih:1970:TDL

[23] Chau Shih and W. J. Karplus. Turbulent diffusion in the lower atmosphere: a hybrid simulation study. Mathematics and Computers in Simulation, 12(3):130–137, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800118.

Gonzalez:1970:OSH

[24] Robert S. Gonzalez. An optimization study on a hybrid computer. Mathematics and Computers in Simulation, 12(3):138–148, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547080012X.

Anonymous:1970:RAc

[25] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 12(3):149–150, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800131.

Anonymous:1970:CBc

[26] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 12(3):151–156, July 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800143.

Anonymous:1970:DRc

[27] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 12(3):157–160, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800155.

Anonymous:1970:Ic

[28] Anonymous. Informations. Mathematics and Computers in Simulation, 12(3):161–167, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800167.

Anonymous:1970:ALc

[29] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 12(3):168–169, July 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800179.

Anonymous:1970:RAT

[30] Anonymous. Recommendations to authors of texts for publication in the Annals of the International Association for Analog Computation. Mathematics and Computers in Simulation, 12(3):170, July 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800180.

Anonymous:1970:PJb

[31] Anonymous. Pages 118–170 (July 1970). Mathematics and Computers in Simulation, 12(3):??, July 1970. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Honnell:1970:LFE

[32] P. M. Honnell, K. J. Borgwald, and R. J. Mulholland. Linear functional entries for the Matric Computor. Mathematics and Computers in Simulation, 12 (4):172–181, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800295.

Kremen:1970:CPI

[33] Jaromir Kremen and Petr Demel. Contribution to the problem of initial conditions of linear dynamical systems at a discontinuous driving function. Mathematics and Computers in Simulation, 12(4):182–191, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800301.

Jehly:1970:SFI

[34] Lynn C. Jehly and Everett D. Alton. Simulation of a flight instrument that indicates and reports true altitude of an aircraft. Mathematics and Computers in Simulation, 12 (4):192–203, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800313.

Anonymous:1970:RAd

[35] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 12(4):204–205, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800325.

Anonymous:1970:CBd

[36] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 12(4):206–211, October 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800337.

Anonymous:1970:DRd

[37] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 12(4):212–216, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800349.

Anonymous:1970:Id

[38] Anonymous. Informations. Mathematics and Computers in Simulation, 12 (4):217–220, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475470800350.

Anonymous:1970:ALd

[39] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 12(4):221–222, October 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800362.

Anonymous:1970:IS

[40] Anonymous. Index — sommaire. Mathematics and Computers in Simulation, 12(4):223–224, October 1970. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475470800374.

Anonymous:1970:PO

[41] Anonymous. Pages 172–224 (October 1970). Mathematics and Computers in Simulation, 12(4):??, October 1970. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Miron:1971:SER

[42] C. Miron and Lelia Peteanu. Simulation exacte d’un retardateur pour fonctions polynomiales. (French) [Exact simulation of a delay for polynomial functions]. Mathematics and Computers in Simulation, 13(1):2–5, January 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800186.

Baert:1971:SNL

[43] D. Baert. A simulated non-linear self-inductance. Mathematics and Computers in Simulation, 13(1):6– 10, January 1971. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800198.

Garcia:1971:AUP

[44] B. Garcia and A. Civit. Analog unit for the performance of ∫_{0}^{t} f(t) dg(t). Mathematics and Computers in Simulation, 13(1):11–17, January 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800204.

Sahni:1971:ACN

[45] R. N. Sahni. Analog computer for non-linear optimal control. Mathematics and Computers in Simulation, 13(1):18–24, January 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800216.

Anonymous:1971:RAa

[46] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 13(1):25–26, January 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800228.

Anonymous:1971:CBa

[47] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 13(1):27–32, January 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847547180023X.

Anonymous:1971:DRa

[48] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 13(1):33–36, January 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800241.

Anonymous:1971:Ia

[49] Anonymous. Informations. Mathematics and Computers in Simulation, 13(1):37–45, January 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800253.

Anonymous:1971:ALa

[50] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 13(1):46–51, January 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800265.

Anonymous:1971:PJa

[51] Anonymous. Pages 2–51 (January 1971). Mathematics and Computers in Simulation, 13(1):??, January 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Barugola:1971:DDP

[52] A. Barugola and J. Y. Gay. Un dispositif de dérivation par échantilonnage pour signaux de très basses fréquences. (French) [A sampling derivation device for very low frequency signals]. Mathematics and Computers in Simulation, 13(2):55–59, April 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800095.

Lauw:1971:SBS

[53] H. K. Lauw. Simulation of a brushless synchronous machine with an analog computer. Mathematics and Computers in Simulation, 13(2):60–75, April 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800101.

Ricci:1971:CSC

[54] Fred J. Ricci. Computer simulation of cybernetic systems. Mathematics and Computers in Simulation, 13(2):76–81, April 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800113.

Midgley:1971:IFP

[55] Denis Midgley. An incremental field plotting method for analogue computers. Mathematics and Computers in Simulation, 13(2):82–84, April 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475471800125.

Anonymous:1971:RAb

[56] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 13(2):85–86, April 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475471800137.

Anonymous:1971:CBb

[57] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 13(2):87–94, April 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800149.

Anonymous:1971:DRb

[58] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 13(2):95–97, April 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800150.

Anonymous:1971:Ib

[59] Anonymous. Informations. Mathematics and Computers in Simulation, 13(2):98–106, April 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800162.

Anonymous:1971:ALb

[60] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 13(2):107–108, April 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800174.

Anonymous:1971:PA

[61] Anonymous. Pages 55–108 (April 1971). Mathematics and Computers in Simulation, 13(2):??, April 1971. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Engle:1971:HDR

[62] Ingrid A. Engle. Hybrid data reduction for inflight analysis of a missile’s warhead adaption kit and fuze. Mathematics and Computers in Simulation, 13(3):111–115, July 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800277.

Ricci:1971:CSP

[63] Fred Ricci. Computer simulation and parameter optimization of phaselocked loops (PLL). Mathematics and Computers in Simulation, 13 (3):116–123, July 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800289.

Tschann:1971:PLG

[64] C. Tschann, V. J. Modi, and A. Soudack. Planar librations of gravity-oriented satellites using analog simulation. Mathematics and Computers in Simulation, 13(3):124–130, July 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800290.

Anonymous:1971:RAc

[65] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 13(3):131–132, July 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800307.

Anonymous:1971:CBc

[66] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 13(3):133–138, July 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800319.

Anonymous:1971:DRc

[67] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 13(3):139–142, July 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800320.

Anonymous:1971:Ic

[68] Anonymous. Informations. Mathematics and Computers in Simulation, 13(3):143–146, July 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800332.

Anonymous:1971:ALc

[69] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 13(3):147–148, July 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800344.

Anonymous:1971:PJb

[70] Anonymous. Pages 111–148 (July 1971). Mathematics and Computers in Simulation, 13(3):??, July 1971. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liebert:1971:DIF

[71] Thomas A. Liebert. Dare II: Fast on-line digital simulation on a small computer.

Mathematics and Computers in Simulation, 13(4):150–155, October 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800010.

Rodriguez-Izquierdo:1971:HGI

[72] G. Rodriguez-Izquierdo and A. Civit. High gain integrated amplifiers stabilize function generators. Mathematics and Computers in Simulation, 13 (4):156–159, October 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800022.

Suske:1971:DVC

[73] Arnost Suske. Le dépouillement des valeurs caractéristiques du mouvement des véhicules sur calculatrice analogique. (French) [Calculation of characteristic values of vehicle movement on an analog calculator]. Mathematics and Computers in Simulation, 13 (4):160–164, October 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800034.

Anonymous:1971:RAd

[74] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 13(4):165, October 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800046.

Anonymous:1971:CBd

[75] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 13(4):166–176, October 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800058.

Anonymous:1971:DRd

[76] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 13(4):177–182, October 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547180006X.

Anonymous:1971:Id

[77] Anonymous. Informations. Mathematics and Computers in Simulation, 13 (4):183–193, October 1971. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800071.

Anonymous:1971:ALd

[78] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 13(4):194–197, October 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475471800083.

Anonymous:1971:PO

[79] Anonymous. Pages 150–197 (October 1971). Mathematics and Computers in Simulation, 13(4):??, October 1971. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Feilmeier:1972:HDG

[80] M. Feilmeier. Die Hybride Darstellung von Grössen in Differentialgleichungen. (German) [The hybrid representation of size in differential equations]. Mathematics and Computers in Simulation, 14(1):2–5, January 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800132.

Soldat:1972:API

[81] Georges Soldat and Guy Mesnard. Automatisation d’un processus itératif convergent de résolution de systèmes linéaires homogènes et paramétriques par un calculateur hybride. (French) [Automation of a convergent iterative process of resolution of homogeneous and parametric linear systems by a hybrid calculator]. Mathematics and Computers in Simulation, 14(1):6–9, January 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800144.

Honnell:1972:DEA

[82] Pierre M. Honnell and Kenneth J. Borgwald. Dynamo-electric amplifiers with Matric Computor feedback control. Mathematics and Computers in Simulation, 14(1):10–16, January 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800156.

Kogan:1972:ACH

[83] B. Ya. Kogan, M. V. Rybashov, and F. M. Babushkin. An analog computer in hybrid loops. Mathematics and Computers in Simulation, 14 (1):17–31, January 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800168.

Anonymous:1972:RAa

[84] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 14(1):32, January 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547280017X.

Anonymous:1972:CBa

[85] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 14(1):33–40, January 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800181.

Anonymous:1972:DRa

[86] Anonymous. Documents recus. Mathematics and Computers in Simulation, 14(1):41–45, January 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800193.

Anonymous:1972:Ia

[87] Anonymous. Informations. Mathematics and Computers in Simulation, 14(1):46–50, January 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547280020X.

Anonymous:1972:AA

[88] Anonymous. Activite de l’association. Mathematics and Computers in Simulation, 14(1):51–52, January 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800211.

Anonymous:1972:PJa

[89] Anonymous. Pages 2–52 (January 1972). Mathematics and Computers in Simulation, 14(1):??, January 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pettersen:1972:GFA

[90] Odd Pettersen and William J. Vetter. A general formulation of analogue computer programming and scaling. Mathematics and Computers in Simulation, 14(2):55–62, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800326.

vanderMast:1972:PRU

[91] C. A. P. G. van der Mast. Pattern recognition using a hybrid computer. Mathematics and Computers in Simulation, 14(2):63–64, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800338.

Malavard:1972:RAC

[92] L. Malavard, G. Renard, and J. J. Rohner. La représentation analogique des champs harmoniques tridimensionnels indefinis et son application au champ électrique cardiaque.

(French) [The analog representation of indefinite three-dimensional harmonic fields and its application to the cardiac electric field]. Mathematics and Computers in Simulation, 14(2):65–69, April 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547280034X.

Buis:1972:HST

[93] J. P. Buis. Hybrid simulation of the transient heat-transfer in a fixed bed with axial conductivity. Mathematics and Computers in Simulation, 14(2):71–80, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800351.

Anonymous:1972:RAb

[94] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 14(2):81–82, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800363.

Anonymous:1972:CBb

[95] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 14(2):83–96, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800375.

Anonymous:1972:DRb

[96] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 14(2):97–99, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800387.

Anonymous:1972:Ib

[97] Anonymous. Informations. Mathematics and Computers in Simulation, 14(2):100–110, April 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800399.

Anonymous:1972:ALa

[98] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 14(2):111–112, April 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800405.

Anonymous:1972:PA

[99] Anonymous. Pages 55–112 (April 1972). Mathematics and Computers in Simulation, 14(2):??, April 1972. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1972:E

[100] Anonymous.  Errata.  Mathemat-

ics and Computers in Simulation, 14 (3):i, July 1972. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800120.

Forestier:1972:UDCa

[101] Christian Forestier and Guy Mesnard. Utilisation du calcul hybride à la détermination de la fonction de partition et des fonctions thermodynamiques d’une particule mobile dans un puits de potentiel. (French) [Use of hybrid calculation to determinate the partition function and the thermodynamic functions of a mobile particle in a potential well]. Mathematics and Computers in Simulation, 14(3):114–126, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800016.

Hien:1972:RES

[102] Vo Hoang Hien and Guy Mesnard.

Résolution de l’équation de Schrödinger à une dimension à l’aide d’un réseau de résistances. (French) [Resolution of the equation of Schrödinger in one dimension using a network of resistances]. Mathematics and Computers in Simulation, 14(3):127–129, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800028.

Bonnemay:1972:SDB

[103] Alain Bonnemay. Stabilité dynamique d’une boucle analogique et précision de la résolution. (French) [Dynamic stability of an analog loop and precision of resolution]. Mathematics and Computers in Simulation, 14 (3):130–135, July 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547280003X.

Iveraen:1972:DSA

[104] J. D. Iveraen and S. Bernstein. Dynamic simulation of an aircraft under the effect of vortex wake turbulence. Mathematics and Computers in Simulation, 14(3):136–144, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800041.

Cocquyt:1972:SMP

[105] G. Cocquyt and G. Vloemans. Simulation of a mechanical pendulum and representation of the corresponding cylindrical phase surface. Mathematics and Computers in Simulation, 14(3):145–148, July 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800053.

Necula:1972:EDC

[106] Maria Ana Necula and Fänicä Spinei. Étude du champ magnétique dans des tôles ferromagnétiques au moyen d’un calculateur analogique. (French) [Study of the magnetic field in ferromagnetic sheets by means of an analog calculator]. Mathematics and Computers in Simulation, 14(3):149–151, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800065.

Anonymous:1972:RAc

[107] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 14(3):152–153, July 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800077.

Anonymous:1972:CBc

[108] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 14(3):154–165, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800089.

Anonymous:1972:DRc

[109] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 14(3):166–173, July 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800090.

Anonymous:1972:Ic

[110] Anonymous. Informations. Mathematics and Computers in Simulation, 14(3):174–178, July 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800107.

Anonymous:1972:ALb

[111] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 14(3):178–179, July 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800119.

Anonymous:1972:PJb

[112] Anonymous. Pages 114–179 (July 1972).

Mathematics and Computers in Simulation, 14(3):??, July 1972. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Forestier:1972:UDCb

[113] Christian Forestier and Guy Mesnard. Utilisation du calcul analogique répétitif à la résolution de l’équation de Schrödinger dans le cas d’un puits de potentiel quelconque. (French) [Use of repetitive analog calculation to the resolution of the equation of Schrödinger in the case of any potential well]. Mathematics and Computers in Simulation, 14(4):183–190, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800223.

Renard:1972:PIP

[114] Guy Renard, Daniel Teil, and Joseph Zarka. Processus itératif pour l’imposition des conditions aux limites en calcul hybride par réseaux. (French) [Iterative process for the imposition of limits in hybrid calculation by networks]. Mathematics and Computers in Simulation, 14(4):191–195, October 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800235.

Jaroslaw:1972:MAS

[115] Pakleza Jaroslaw, Lobodzinski Wojciech, and Osorio Angel. Une méthode analogique de simulation des champs thermiques stationnaires non linéaires en presence de sources internes de chaleur. (French) [An analog simulation of nonlinear stationary thermal fields in the presence of internal heat sources]. Mathematics and Computers in Simulation, 14(4):196–206, October 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800247.

Roux:1972:GSP

[116] P. Roux. Générateur de séquence programmable. (French) [Programmable sequence generator]. Mathematics and Computers in Simulation, 14(4): 207–209, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800259.

Anonymous:1972:RAd

[117] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 14(4):210, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800260.

Anonymous:1972:CBd

[118] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 14(4):211–222, October 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800272.

Anonymous:1972:DRd

[119] Anonymous. Documents recus. Mathematics and Computers in Simulation, 14(4):223–226, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475472800284.

Anonymous:1972:Id

[120] Anonymous. Informations. Mathematics and Computers in Simulation, 14 (4):226–232, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800296.

Anonymous:1972:ALc

[121] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 14(4):230–231, October 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800302.

Anonymous:1972:Ie

[122] Anonymous. Index 1971–72. Mathematics and Computers in Simulation, 14(4):232, October 1972. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475472800314.

Anonymous:1972:PO

[123] Anonymous. Pages 183–232 (October 1972). Mathematics and Computers in Simulation, 14(4):??, October 1972. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Coeurdeuil:1973:RCE

[124] M. M. Gilles Coeurdeuil and Romane Mezencev. Résolution de certaines équations intégrales et intégrodifférentialles sur calculateur hybride. (French) [Resolution of certain integral and integro-differential equations on a hybrid calculator]. Mathematics and Computers in Simulation, 15(1):3–7, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800011.

vandeVen:1973:MIA

[125] H. H. van de Ven, A. P. Gisling, and P. Eykhoff. The measurement of the impedance by automatic adjustment of a physical model. Mathematics and Computers in Simulation, 15(1):8–15, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800023.

Churassy:1973:SEE

[126] Serge Churassy and Guy Mesnard. Simulation des états électroniques dans les solides au moyen d’un calculateur analogique à réseau de résistances. (French) [Simulation of electronic states in solids by means of an analog calculator with a network of resistors]. Mathematics and Computers in Simulation, 15(1):16–18, January 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800035.

Sharafedin:1973:SEA

[127] Ahmad Sharafedin. Solution de l’équation algébrique par une méthode analogique. Application aux équations algébriques à coefficients paramétriques. (French) [Algebraic equation solution by an analog method. Application to algebraic equations with parametric coefficients]. Mathematics and Computers in Simulation, 15 (1):19–24, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800047.

Anonymous:1973:RAa

[128] Anonymous. Resumes d’articles. Mathematics and Computers in Simulation, 15(1):25–26, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800059.

Anonymous:1973:CBa

[129] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 15(1):27–33, January 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800060.

Anonymous:1973:DRa

[130] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 15(1):34–39, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800072.

Anonymous:1973:Ia

[131] Anonymous. Informations. Mathematics and Computers in Simulation, 15(1):40–44, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800084.

Anonymous:1973:AAa

[132] Anonymous. Activite de l’association.

Mathematics and Computers in Simulation, 15(1):45–46, January 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800096.

Anonymous:1973:PJa

[133] Anonymous. Pages 3–46 (January 1973).

Mathematics and Computers in Simulation, 15(1):??, January 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Khalil:1973:HCS

[134] T. M. Khalil, P. E. Valisalo, and A. Copsey. Hybrid computer solution of queueing systems. Mathematics and Computers in Simulation, 15(2):49–53, April 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800102.

Rozanski:1973:RME

[135] Lech Różański. Réalisation des modèles électriques pour des systèmes thermiques basée sur le Théorème de Thévenin. (French) [Realization of electrical models for thermal systems based on Thévenin’s theorem]. Mathematics and Computers in Simulation, 15(2):54–61, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800114.

Caussade:1973:PCH

[136] B. Caussade, J. Piquemal, and C. Zarcone. Principes d’un calculateur hybride associant un calculateur numérique à un réseau de résistances. (French) [Principles of a hybrid calculator combining a numerical calculator with a network of resistances]. Mathematics and Computers in Simulation, 15(2):62–64, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800126.

Hien:1973:SMA

[137] Vo Hoang Hien and Guy Mesnard. Stabilité des montages à amplificateur opérationnel. Application aux résistances négatives. (French) [stability of the operational amplifier assemblies. Application to negative resistances]. Mathematics and Computers in Simulation, 15(2):65–71, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800138.

Anonymous:1973:RAb

[138] Anonymous. Resumés d’articles. Mathematics and Computers in Simulation, 15(2):72–73, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547380014X.

Anonymous:1973:CBb

[139] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 15(2):74–77, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800151.

Anonymous:1973:DRb

[140] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 15(2):78–80, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800163.

Anonymous:1973:Ib

[141] Anonymous. Informations. Mathematics and Computers in Simulation, 15(2):81–84, April 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800175.

Anonymous:1973:AAb

[142] Anonymous. Activite de l’association.

Mathematics and Computers in Simulation, 15(2):85, April 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800187.

Anonymous:1973:PA

[143] Anonymous. Pages 49–85 (April 1973).

Mathematics and Computers in Simulation, 15(2):??, April 1973. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Gyergyek:1973:CES

[144] L. Gyergyek and R. Babic. Contribution to the estimation of the speech message from the corrupted signal with noise. Mathematics and Computers in Simulation, 15(3):89–91, July 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800278.

Mesnard:1973:SCR

[145] Guy Mesnard, Vo Hoang Hien, and Jean Vidal. Stabilite des calculateurs a reseaux comportant des resistances negatives. Mathematics and Computers in Simulation, 15(3):92–96, July 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547380028X.

Hien:1973:PSR

[146] Vo Hoang Hien, Guy Mesnard, and Lionel Heitz. Procédures de stabilisation des réseaux comportant des résistances négatives. (French) [network stabilization procedures comprising negative resistances]. Mathematics and Computers in Simulation, 15(3):97–103, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800291.

Anonymous:1973:E

[147] Anonymous.  Erratum.  Mathe-

matics and Computers in Simulation, 15(3):103, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800308.

Genin:1973:EAE

[148] J. Genin and R. Genin. Etude analogique des equations. Mathematics and Computers in Simulation, 15(3):104–114, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547380031X.

Bastin:1973:PCC

[149] Henri Bastin and Monique Delchambre. Procedure De Calcul Des Coefficients Des Series De Froebenius Solutions Des Equations Fuchsiennes. Mathematics and Computers in Simulation, 15(3):115–119, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800321.

Anonymous:1973:RAc

[150] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 15(3):120, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800333.

Anonymous:1973:CBc

[151] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 15(3):121–128, July 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800345.

Anonymous:1973:DRc

[152] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 15(3):129–133, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800357.

Anonymous:1973:Ic

[153] Anonymous. Informations. Mathematics and Computers in Simulation, 15(3):134–140, July 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800369.

Anonymous:1973:AAc

[154] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 15(3):141–142, July 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800370.

Anonymous:1973:PJb

[155] Anonymous. Pages 89–142 (July 1973). Mathematics and Computers in Simulation, 15(3):??, July 1973. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Demoment:1973:SCA

[156] Guy Demoment, Pierre Peronneau, Raymond Duperdu, and Jacques Hinglais. Simulation sur calculateur analogique du fonctionnement mécanique du système cardiovasculaire. (French) [simulation on an analog calculator of the mechanical functioning of the cardiovascular system]. Mathematics and Computers in Simulation, 15 (4):145–158, October 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800199.

Hogg:1973:SAS

[157] B. W. Hogg and D. Williams. Simulation of an asynchronized synchronous machine. Mathematics and Computers in Simulation, 15(4):159– 167, October 1973. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800205.

Ricci:1973:SAA

[158] Ferdinand J. Ricci. Systems analysis applied to a hybrid computer simulation of a missile reentering the atmosphere. Mathematics and Computers in Simulation, 15(4):168–174, October 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800217.

Anonymous:1973:RAd

[159] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 15(4):175–176, October 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800229.

Anonymous:1973:CBd

[160] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 15(4):177–186, October 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475473800230.

Anonymous:1973:DRd

[161] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 15(4):187–189, October 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800242.

Anonymous:1973:Id

[162] Anonymous. Informations. Mathematics and Computers in Simulation, 15 (4):190–193, October 1973. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800254.

Anonymous:1973:AL

[163] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 15(4):194–197, October 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475473800266.

Anonymous:1973:PO

[164] Anonymous. Pages 145–197 (October 1973). Mathematics and Computers in Simulation, 15(4):??, October 1973. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1974:E

[165] R. Vichnevetsky. Editorial. Mathematics and Computers in Simulation, 16(1):1–2, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800017.

Vichnevetsky:1974:PCC

[166] Robert Vichnevetsky. Physical criteria in computer methods for partial differential equations. Mathematics and Computers in Simulation, 16(1):3–16, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800029.

Nilsen:1974:CSS

[167] Ragnar N. Nilsen and Walter J.

Karplus. Continuous-system simulation languages: a state-of-the-art survey. Mathematics and Computers in Simulation, 16(1):17–25, January 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800030.

Rideout:1974:CSS

[168] Vincent C. Rideout and Maurice F.

Snyder. Computer simulation study of the cardiovascular and related physiological systems. Mathematics and Computers in Simulation, 16 (1):26–34, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800042.

Clemence:1974:DCA

[169] C. R. Clemence. Digital control of an analog marketing model. Mathematics and Computers in Simulation, 16(1):35–39, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800054.

Costanti:1974:MPF

[170] F. J. Costanti and J. A. R. Santos. A model for the Purkinje fibers action potential, using a Hybrid Computer. Mathematics and Computers in Simulation, 16(1):40–43, January 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800066.

Anonymous:1974:Ba

[171] Anonymous. Bibliographie. Mathematics and Computers in Simulation, 16(1):44–50, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800078.

Anonymous:1974:PJa

[172] Anonymous. Pages 1–50 (January 1974). Mathematics and Computers in Simulation, 16(1):??, January 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rideout:1974:E

[173] V. C. Rideout. Editorial. Mathematics and Computers in Simulation, 16(2):1, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847547480008X.

Korn:1974:RCS

[174] Granino A. Korn. Recent computer system developments and continuous system simulation. Mathematics and Computers in Simulation, 16(2):2–11, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800091.

Hughes:1974:HOI

[175] J. R. Dixon Hughes, V. R. Sastry, and D. G. Wong. Higher order incremental integrators for DDA type computer structures. Mathematics and Computers in Simulation, 16(2):12–19, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800108.

Vichnevetsky:1974:CAH

[176] R. Vichnevetsky and A. W. Tomalesky.

Computer algorithm for hyperbolic equations: With application to the simulation of electrical-transmission lines. Mathematics and Computers in Simulation, 16(2):20–25, April 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547480011X.

Preucil:1974:HRI

[177] M. Preucil. Hybrid reconstructors — Interpolators. Mathematics and Computers in Simulation, 16(2):26–37, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800121.

Barker:1974:EAP

[178] Derek C. Barker. Electrical analogue for a partial differential equation. Mathematics and Computers in Simulation, 16(2):38–45, April 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800133.

Galletti:1974:SDS

[179] Carlo Galletti, Aleramo Lucifredi, and Aldo Rossi. On the simulation of dynamic systems and their graphic display by hybrid computer. Mathematics and Computers in Simulation, 16(2):46–51, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800145.

Anonymous:1974:Bb

[180] Anonymous. Bibliographie. Mathematics and Computers in Simulation, 16(2):52–58, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800157.

Anonymous:1974:PA

[181] Anonymous. Pages 1–58 (April 1974). Mathematics and Computers in Simulation, 16(2):??, April 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vansteenkiste:1974:E

[182] C. C. Vansteenkiste. Editorial. Mathematics and Computers in Simulation, 16(3):1, July 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847547480025X.

Tam:1974:PIM

[183] Wing Cheung Tam and Walter J.

Karplus. Parameter identification in models of hydrologic systems using the digital simulation language PDEL. Mathematics and Computers in Simulation, 16(3):2–8, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800261.

Bishop:1974:AHA

[184] Kenneth A. Bishop. An application of the hybrid alternating direction implicit procedure to the simulation of an unconfined aquifer. Mathematics and Computers in Simulation, 16(3):9–14, July 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800273.

Davidson:1974:SOC

[185] Burton Davidson and A. K. Mendiratta. Simulation of the optimal combination of waste treatment and spatially distributed discharge of effluent using Pontryagin’s minimum principle. Mathematics and Computers in Simulation, 16(3):15–30, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800285.

Vansteenkiste:1974:IMT

[186] G. C. Vansteenkiste and F. De Schutter. Implementation of the modal technique to diffusion processes. Mathematics and Computers in Simulation, 16(3):31–34, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800297.

Renard:1974:ASD

[187] G. Renard, M. Aucouturier, and J. P. Yaher. Analog study of the draw down of the groundwater table: the need for an interface. Mathematics and Computers in Simulation, 16(3):35–45, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800303.

Ung:1974:AHC

[188] Man T. Ung, George A. Bekey, and Michael J. Merritt. Analysis of hybrid computer interface errors on the solution of partial differential equations. Mathematics and Computers in Simulation, 16(3):47–58, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800315.

Anonymous:1974:Bc

[189] Anonymous. Bibliography. Mathematics and Computers in Simulation, 16(3):59–74, July 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475474800327.

Anonymous:1974:PJb

[190] Anonymous. Pages 1–74 (July 1974). Mathematics and Computers in Simulation, 16(3):??, July 1974. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cennamo:1974:OQF

[191] Felice Cennamo and Pier Luigi Rocca.

Optimization of quadrature formulas for a modular digital differential analyzer. Mathematics and Computers in Simulation, 16(4):1–9, October 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800169.

Burnett:1974:MMI

[192] Robert A. Burnett, Paul J. Dionne, Eugene A. Eschbach, William A. Reardon, L. Charles Miller, and David A. Summers. A man-model interactive computer system for analysis of complex societal problems. Mathematics and Computers in Simulation, 16 (4):10–13, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800170.

Dekker:1974:APN

[193] L. Dekker and E. H. J. Kerckhoffs. Automatized programming of nonlinear differential systems in a hybrid computer. Hybrid simulation of a world model. Mathematics and Computers in Simulation, 16(4):14–23, October 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800182.

Bosgra:1974:FAS

[194] Okko H. Bosgra and J. Paul Buis.

Fundamental approach to the selection of implementation parameters in the hybrid serial solution of partial differential equations. Mathematics and Computers in Simulation, 16 (4):24–34, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800194.

Kohne:1974:MSM

[195] M. Köhne and M. Zeitz. Modal simulation method modifications and implementations. Mathematics and Computers in Simulation, 16(4):35– 38, October 1974. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800200.

Barczak:1974:SPM

[196] C. L. Barczak. Simulation of a period measurer for low frequencies. Mathematics and Computers in Simulation, 16(4):39–40, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800212.

Hojberg:1974:DPH

[197] K. Søe Højberg and Jens V. Olsen. Digital potentiometer for the hybrid computer EAI 680–PDP 8/I. Mathematics and Computers in Simulation, 16(4):41–43, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800224.

Anonymous:1974:CB

[198] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 16(4):44–56, October 1974. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800236.

Anonymous:1974:IS

[199] Anonymous. Index — Sommaire. Mathematics and Computers in Simulation, 16(4):57–58, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475474800248.

Anonymous:1974:PO

[200] Anonymous. Pages 1–58 (October 1974).

Mathematics and Computers in Simulation, 16(4):??, October 1974. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1975:MSI

[201] R. Vichnevetsky. Modeling, simulation and interdisciplinary research. Mathematics and Computers in Simulation, 17(1):1–2, January 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800401.

Bellini:1975:NAS

[202] Armando Bellini, Claudio Cerri, and Alessandro De Carli. A new approach to the simulation of static converter drives. Mathematics and Computers in Simulation, 17(1):3–7, January 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800413.

Piceni:1975:UHC

[203] Hans A. L. Piceni and Pieter Eykhoff. The use of hybrid computers for system-parameter estimation. Mathematics and Computers in Simulation, 17(1):8–22, January 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800425.

Rideout:1975:PEA

[204] V. C. Rideout and J. E. W. Beneken. Parameter estimation applied to physiological systems. Mathematics and Computers in Simulation, 17(1):23– 36, January 1975. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800437.

Bellini:1975:SPC

[205] Armando Bellini and Alessandro De Carli. Simulation of phase-controlled induction motors. Mathematics and Computers in Simulation, 17(1):37– 43, January 1975. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800449.

Anonymous:1975:CBa

[206] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 17(1):44–56, January 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800450.

Anonymous:1975:PJa

[207] Anonymous. Pages 1–56 (January 1975).

Mathematics and Computers in Simulation, 17(1):??, January 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1975:E

[208] Robert Vichnevetsky. Editorial. Mathematics and Computers in Simulation, 17(2):57, April 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800139.

Hien:1975:ICL

[209] Vo hoang Hien and Guy Mesnard. Introduction des conditions aux limites dans les calculateurs à réseau électrique à l’aide d’éléments pathologiques. (French) [introduction of boundary conditions in electric network calculators using pathological elements]. Mathematics and Computers in Simulation, 17(2):61–65, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800140.

Mesnard:1975:CRR

[210] Guy Mesnard and Vo-Hoang Hien. Calculateurs à réseaux de résistances fonctionnant en régime alternatif. (French) [Resistance network calculators operating in an alternative regime]. Mathematics and Computers in Simulation, 17(2):66–68, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800152.

Govindarao:1975:ASI

[211] V. M. H. Govindarao. Analog simulation of an isothermal semibatch bubble-column slurry reactor. Mathematics and Computers in Simulation, 17(2):69–78, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800164.

Tarela:1975:MLPa

[212] J. M. Tarela, J. M. Perez, and V. Aleixandre. Minimization of lattice polynomials on piecewise linear functions (Part I). Mathematics and Computers in Simulation, 17(2):79–85, April 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800176.

Birkhoff:1975:NSH

[213] Garrett Birkhoff and Vassilios A. Dougalis. Numerical solution of hydrodynamic problems. Mathematics and Computers in Simulation, 17(2):88–94, April 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800188.

Swartz:1975:CCC

[214] Blair Swartz. Comparing certain classes of difference and finite element methods for a hyperbolic problem. Mathematics and Computers in Simulation, 17(2):95–97, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547580019X.

Houstis:1975:DES

[215] E. N. Houstis, R. E. Lynch, T. S.

Papatheodorou, and J. R. Rice. Development, evaluation and selection of methods for elliptic partial differential equations. Mathematics and Computers in Simulation, 17(2):98– 103, April 1975. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800206.

Petrov:1975:DDT

[216] Andrey A. Petrov. A differentialdifference technique for the hybrid computer solution of parabolic partial differential equations. Mathematics and Computers in Simulation, 17 (2):104–109, April 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800218.

Humeau:1975:HSC

[217] J. P. Humeau and M. Grolleau. Hybrid simulation of a control problem for a distributed parameter system. Mathematics and Computers in Simulation, 17(2):110–115, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547580022X.

Shaug:1975:SFP

[218] James C. Shaug and John C. Bruch. Seepage flow problems in layered soils solved by the finite element method. Mathematics and Computers in Simulation, 17(2):116–120, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800231.

Tarela:1975:MLPb

[219] J. M. Tarela, L. A. Bailon, and E. Sanz G. Minimization of lattice polynomials on piecewise linear functions (Part II). Mathematics and Computers in Simulation, 17(2):121–127, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800243.

Bernard:1975:CHR

[220] Caussade Bernard, Renard Guy, Zarcone César, and Bourrel Jean. Un calculateur hybride à réseau d’impédances. (French) [A hybrid calculator with impedance network]. Mathematics and Computers in Simulation, 17 (2):128–136, April 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800255.

Kerneuzet:1975:MSF

[221] A. Kerneuzet and R. Genin. Sur une méthode de simulation des fonctions retardées. (French) [on a simulation method for delayed functions]. Mathematics and Computers in Simulation, 17(2):137–141, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800267.

Anonymous:1975:CBb

[222] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 17(2):142–152, April 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800279.

Anonymous:1975:PA

[223] Anonymous. Pages 57–152 (April 1975).

Mathematics and Computers in Simulation, 17(2):??, April 1975. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mesnard:1975:EPV

[224] Guy Mesnard and Vo hoang Hien.

Études des problèmes de valeurs propres avec des calculateurs à réseaux électriques emploi d’éléments pathologiques. Mathematics and Computers in Simulation, 17(3):153–156, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800462.

Znamirowski:1975:SPD

[225] L. Znamirowski and O. A. Palusinski. Scaling problems in dynamic system simulation. Mathematics and Computers in Simulation, 17 (3):157–162, July 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800474.

Overington:1975:DSC

[226] William J. G. Overington. A design study for a Chebyshev polynomial unit for an analogue computer. Mathematics and Computers in Simulation, 17(3):163–170, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800486.

Ruppel:1975:ISM

[227] W. Ruppel and M. Zeitz. Implicit simulation method for distributed parameter. Mathematics and Computers in Simulation, 17(3):171–177, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800498.

Vichnevetsky:1975:FAF

[228] R. Vichnevetsky and F. De Schutter. A frequency analysis of finite difference and finite element methods for initial value problems. Mathematics and Computers in Simulation, 17(3):179–185, July 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800504.

Kincaid:1975:DCP

[229] David R. Kincaid and David M. Young. The development of a computer package for solving a class of partial differential equations by iterative methods. Mathematics and Computers in Simulation, 17(3):186–191, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800516.

Gary:1975:CSF

[230] John Gary. Comparison of some fourth order difference schemes for hyperbolic problems. Mathematics and Computers in Simulation, 17 (3):192–194, July 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800528.

Carver:1975:FSA

[231] M. B. Carver. The FORSIM system for automated solution of sets of implicitly coupled partial differential equations. Mathematics and Computers in Simulation, 17(3):195–202, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547580053X.

OBrien:1975:HCA

[232] Joseph F. O’Brien and Billy L. Edge.

A hybrid computer approach to the solution of water quality in an estuary. Mathematics and Computers in Simulation, 17(3):203–210, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800541.

Bracho:1975:TFC

[233] S. Bracho and A. Buron. Transfer function can be controlled and generated by means of stabilized feedback multiplier. Mathematics and Computers in Simulation, 17(3):211–214, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800553.

Anonymous:1975:CBc

[234] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 17(3):215–230, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800565.

Anonymous:1975:PJb

[235] Anonymous. Pages 153–230, 1-19 (July 1975). Mathematics and Computers in Simulation, 17(3):??, July 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Elzas:1975:E

[236] M. S. Elzas. Editorial. Mathematics and Computers in Simulation, 17(4):231, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800012.

Dekker:1975:AHS

[237] L. Dekker. Applicability of hybrid simulation. Mathematics and Computers in Simulation, 17(4):233– 243, October 1975. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800024.

Elzas:1975:LHD

[238] M. S. Elzas and D. Kettenis. Languages for hybrid/digital simulation.

Mathematics and Computers in Simulation, 17(4):245–252, October 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800036.

Timman:1975:PDA

[239] R. Timman, L. Dekker, M. S. Elzas, H. le Grand, J. H. Kruisinga, G. C. Vansteenkiste, and E. J. H. Kerckhoffs. Panel discussion on the applicability of simulation tools. Mathematics and Computers in Simulation, 17 (4):253–259, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800048.

Hay:1975:SLI

[240] J. L. Hay, J. G. Pearce, and M. D.

Narotam. Simulation language implementation on minicomputers. Mathematics and Computers in Simulation, 17(4):260–269, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847547580005X.

Heppner:1975:DDM

[241] Dierk Heppner. DMSL — a descriptive macro simulation language. Mathematics and Computers in Simulation, 17(4):270–276, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800061.

Vajk:1975:PSC

[242] I. Vajk, L. Keviczky, and Cs. Bányász. Procal in the service of control engineering. Mathematics and Computers in Simulation, 17(4): 277–282, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800073.

Mawson:1975:CSS

[243] J. B. Mawson. A continuous system simulation language for an advanced hybrid computing system (Ahcssl). Mathematics and Computers in Simulation, 17(4):283–293, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800085.

Landauer:1975:PGS

[244] J. Paul Landauer. Program generation system for modern hybrid computers. Mathematics and Computers in Simulation, 17(4):294–299, October 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800097.

Bourdeau:1975:UOS

[245] Ch. Bourdeau and J. P. Courtiat. A user-oriented synchronisation system for a hybrid problem. Mathematics and Computers in Simulation, 17 (4):300–302, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475475800103.

Crosbie:1975:ICS

[246] R. E. Crosbie. Interaction in continuoussystem simulation languages. Mathematics and Computers in Simulation, 17(4):303–306, October 1975. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800115.

Anonymous:1975:CBd

[247] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 17(4):309–320, October 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475475800127.

Anonymous:1975:PO

[248] Anonymous. Pages 231–320 (October 1975). Mathematics and Computers in Simulation, 17(4):??, October 1975. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1976:E

[249] Robert Vichnevetsky. Editorial. Mathematics and Computers in Simulation, 18(1):1, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900252.

Bassano:1976:IPH

[250] J. C. Bassano, Y. Lennon, and J. Vignes. An identification of parameters on a hybrid computer. Mathematics and Computers in Simulation, 18(1):3–7, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900264.

Ahmad:1976:CTA

[251] A. I. Sheikh Ahmad. Computer techniques for airfoil design. Mathematics and Computers in Simulation, 18 (1):9–17, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900276.

Pauc:1976:SCS

[252] Gilbert Pauc and Guy Mesnard. Simulation de la circulation systémique du sang par des procédés électriques et analogiques. (French) [systemic blood circulation simulation by electric and analog processes]. Mathematics and Computers in Simulation, 18(1):18–30, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900288.

Robert:1976:MEN

[253] Jean Robert and L. P. A. Robichaud. Mèthode d’ètude numèrique des champs physiques. Application à la dètermination de l’effet Hall dans des plaques rectangulaires. (French) [method of numerical study of physical fields. Application to the determination of the Hall effect in rectangular plates]. Mathematics and Computers in Simulation, 18(1):31–38, January 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547690029X.

Korn:1976:UFH

[254] Granino A. Korn and Olgierd A.

Palusinski. Ultra-fast hybrid analogdigital computation with convenient simulation-language software. Mathematics and Computers in Simulation, 18(1):39–50, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900306.

Kobayashi:1976:SSG

[255] Yasuhiro Kobayashi, Masaaki Ohkita, Michio Inoue, and Masao Nakamura. Separate and/or simultaneous generations of non-linear functions inherent in non-ohmic resistivities. Mathematics and Computers in Simulation, 18(1):51–61, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900318.

Overington:1976:DSI

[256] William J. G. Overington. A design study for an integer order Bessel function of the first kind function generator for an analogue computer. Mathematics and Computers in Simulation, 18(1):63–64, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547690032X.

Anonymous:1976:BCBa

[257] Anonymous. Bibliography chronique bibliographique. Mathematics and Computers in Simulation, 18(1):65– 75, January 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900331.

Anonymous:1976:EB

[258] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 18(1):ifc, January 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900240.

Anonymous:1976:PJa

[259] Anonymous. Pages 1–75 (January 1976). Mathematics and Computers in Simulation, 18(1):??, January 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Saphier:1976:SSG

[260] D. Saphier. Simulation of a steam generator leak detection experiment in the CCTL. Mathematics and Computers in Simulation, 18(2):77– 88, April 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900173.

Yuen:1976:SPC

[261] C. K. Yuen. Some practical considerations concerning piecewise linear approximations. Mathematics and Computers in Simulation, 18(2): 89–91, April 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900185.

Jamshidi:1976:OSD

[262] M. Jamshidi. Optimization of some dynamic industrial control processes by analog simulation. Mathematics and Computers in Simulation, 18 (2):93–100, April 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900197.

Sauvet:1976:CRS

[263] Michel Sauvet and Guy Mesnard. Conception et réalisation d’un simulateur de l’activité électrique cardiaque. (French) [design and realization of a cardiac electric activity simulator]. Mathematics and Computers in Simulation, 18(2):101–107, April 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900203.

Janac:1976:ICI

[264] Karel Janac. Initial condition insertion circuit for hybrid multiplexing systems. Mathematics and Computers in Simulation, 18(2):109–114, April 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900215.

Kobayashi:1976:FPA

[265] Yasuhiro Kobayashi, Masaaki Ohkita, and Michio Inoue. Fractional power approximation and its generation. Mathematics and Computers in Simulation, 18(2):115–122, April 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900227.

Anonymous:1976:B

[266] Anonymous. Bibliography. Mathematics and Computers in Simulation, 18(2):125–138, April 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900239.

Anonymous:1976:PA

[267] Anonymous. Pages 77–138 (April 1976). Mathematics and Computers in Simulation, 18(2):??, April 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

El-Sherif:1976:EAH

[268] A. K. El-Sherif, A. A. Kamal, and A. Y. Bilal. Error analysis in hybrid computer simulation of dynamic systems. Mathematics and Computers in Simulation, 18(3):141–147, July 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900446.

Vichnevetsky:1976:CDF

[269] R. Vichnevetsky. Comments on the derivation of finite-element approximations from conservation principles. Mathematics and Computers in Simulation, 18(3):149–152, July 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900458.

Rochegude:1976:AEU

[270] Patrice Rochegude and Guy Mesnard.

Analogies électromécaniques utilisant des circuits fonctionnant en régime sinusoidal. (French) [Electromechanical analogies using circuits operating in a sinusoidal regime]. Mathematics and Computers in Simulation, 18(3):153–157, July 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547690046X.

Bradley:1976:PIS

[271] J. H. S. Bradley. Precision implicit second order differences for numerical weather simulation. Mathematics and Computers in Simulation, 18(3):159–164, July 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900471.

Fan:1976:SQN

[272] Walter Fan. Simulation of queueing network with time varying arrival rates. Mathematics and Computers in Simulation, 18(3):165–170, July 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900483.

Borsu:1976:SMB

[273] Mme A. Borsu and Ir. L. Gelders.

A simulation model for a bottleneck situation on a river. Mathematics and Computers in Simulation, 18(3):171–176, July 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900495.

Anonymous:1976:E

[274] Anonymous.  Erratum.  Mathe-

matics and Computers in Simulation, 18(3):176, July 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900501.

Caussade:1976:SHP

[275] Bernard Henri Caussade. Simulation hybride de phenomènes de diffusion en régime non linéaire. (French) [Hybrid simulation of diffusion phenomena in a nonlinear regime]. Mathematics and Computers in Simulation, 18(3):177–185, July 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900513.

Overington:1976:CTA

[276] William J. G. Overington. Catastrophe theory and the analogue computer. Mathematics and Computers in Simulation, 18(3):187–188, July 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475476900525.

Anonymous:1976:BCBb

[277] Anonymous. Bibliography chronique bibliographique. Mathematics and Computers in Simulation, 18(3):189– 199, July 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900537.

Anonymous:1976:PJb

[278] Anonymous. Pages 141–199 (July 1976).

Mathematics and Computers in Simulation, 18(3):??, July 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Day:1976:HCI

[279] J. D. Day. Hybrid computer inversion of the Prandtl–Meyer function. Mathematics and Computers in Simulation, 18(4):201–203, October 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547690001X.

Gaubert:1976:SAN

[280] Par G. Gaubert and A. Guillois.

Simulation analogique et numérique par réseaux à éléments variables de l’équation de la diffusion. (French) [Analog and digital simulation by variable element networks of the diffusion equation]. Mathematics and Computers in Simulation, 18(4):205– 216, October 1976. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900021.

Mohapatra:1976:SGR

[281] Pratap K. J. Mohapatra. Some generalised results on exponential delays. Mathematics and Computers in Simulation, 18(4):217–221, October 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900033.

El-Zorkany:1976:HMS

[282] H. I. El-Zorkany and R. Balasubramanian. A hybrid method for the solution of first order partial differential equations arising in non-linear estimation. Mathematics and Computers in Simulation, 18(4):223–230, October 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900045.

Osorio:1976:MND

[283] A. Osorio. Méthodes numériques aux différences finies application aux equations de diffusion non linéaries. (French) [Numerical methods with finite differences: application to nonlinear diffusion equations]. Mathematics and Computers in Simulation, 18 (4):231–238, October 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900057.

Subramanian:1976:MTM

[284] R. Subramanian, D. N. P. Murthy, and T. C. Smith. A method for teaching model-building. Mathematics and Computers in Simulation, 18 (4):239–244, October 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900069.

Kalyayev:1976:HDA

[285] A. V. Kalyayev. Homogeneous digital analogous structures with programmable cojxutation. Mathematics and Computers in Simulation, 18 (4):245–255, October 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900070.

VanRemoortere:1976:BCB

[286] P. Van Remoortere. Bibliography Chronique bibliographique. Mathematics and Computers in Simulation, 18 (4):257–269, October 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900082.

Anonymous:1976:IS

[287] Anonymous. Index — sommaire. Mathematics and Computers in Simulation, 18(4):270–272, October 1976. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475476900094.

Anonymous:1976:PO

[288] Anonymous. Pages 201–272 (October 1976). Mathematics and Computers in Simulation, 18(4):??, October 1976. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1977:E

[289] R. Vichnevetsky. Editorial. Mathematics and Computers in Simulation, 19(1):1, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900337.

Karplus:1977:SMM

[290] Walter J. Karplus. The spectrum of mathematical modeling and systems simulation. Mathematics and Computers in Simulation, 19(1):3– 10, March 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900349.

Roberts:1977:UAC

[291] A. P. Roberts and M. W. A. Smith. Use of an analogue computer in the application of Kalman filter methods of system identification in the presence of noise. Mathematics and Computers in Simulation, 19(1):11–22, March 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900350.

Paker:1977:AMN

[292] Yakup Paker. Application of microprocessor networks for the solution of diffusion equation. Mathematics and Computers in Simulation, 19(1):23–27, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900362.

Fremond:1977:MEF

[293] M. Fremond. La méthode des éleménts finis — application aux équations de la diffusion linéaires et non linéaires. (French) [The finite-element method — application to linear and nonlinear diffusion]. Mathematics and Computers in Simulation, 19(1):29– 37, March 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900374.

Blackstone:1977:CSS

[294] E. H. Blackstone, A. K. Gupta, and V. C. Rideout. Cardiovascular simulation study of infants with transposition of the great arteries after surgical correction. Mathematics and Computers in Simulation, 19 (1):39–50, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900386.

Miller:1977:ERT

[295] A. J. Miller, A. W. Brown, and P. Mars. Error reduction techniques in digital stochastic computers. Mathematics and Computers in Simulation, 19(1):51–56, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900398.

Tang:1977:SDQ

[296] I. Tang. Simultaneous determination of quadratic factors by optimization methods. Mathematics and Computers in Simulation, 19(1):57– 59, March 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900404.

Anonymous:1977:Ba

[297] Anonymous. Bibliography. Mathematics and Computers in Simulation, 19(1):61–74, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900416.

Anonymous:1977:EB

[298] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 19(1):ifc, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900325.

Anonymous:1977:PM

[299] Anonymous. Pages 1–74 (March 1977). Mathematics and Computers in Simulation, 19(1):??, March 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Birolini:1977:HSSa

[300] A. Birolini. Hardware simulation of semi-Markov and related processes: Part I: a versatile generator. Mathematics and Computers in Simulation, 19(2):75–97, June 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900015.

Descusse:1977:DOD

[301] J. Descusse, D. Sarlat, and R. Mezencev. Design of an optimal decoupling control algorithm: (Hybrid simulation test, on-line results). Mathematics and Computers in Simulation, 19 (2):99–112, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900027.

Sinha:1977:LSI

[302] N. K. Sinha and M. Y. Tang. Online system identification in real time using a minicomputer. Mathematics and Computers in Simulation, 19 (2):113–120, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900039.

Hsu:1977:TMV

[303] C. T. Hsu and H. Tesfamariam. Turbulent modeling of a vortex boundary layer flow. Mathematics and Computers in Simulation, 19(2):121– 131, June 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900040.

Mitchell:1977:UCS

[304] Edward E. L. Mitchell. Using a continuous simulation language (ACSL) to model and control a hybrid simulation. Mathematics and Computers in Simulation, 19(2):133–140, June 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900052.

Kobayashi:1977:GTV

[305] Yasuhiro Kobayashi, Masaaki Ohkita, and Michio Inoue. Generation of two-variable functions based on the polynomial approximation. Mathematics and Computers in Simulation, 19(2):141–149, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900064.

Ram:1977:MMI

[306] Giora Ram. M.I.S. — the Meteosat Image Simulator. Mathematics and Computers in Simulation, 19 (2):151–158, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900076.

Vichnevetsky:1977:MSA

[307] Robert Vichnevetsky. Mean squared accuracy of numerical approximations of partial differential equations. Mathematics and Computers in Simulation, 19(2):159–168, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900088.

VanRemoortere:1977:Ba

[308] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 19(2):169–182, June 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847547790009X.

Anonymous:1977:PJ

[309] Anonymous. Pages 75–182 (June 1977). Mathematics and Computers in Simulation, 19(2):??, June 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Birolini:1977:HSSb

[310] A. Birolini. Hardware simulation of semi-Markov and related processes: Part II: Simulation of noise pulses on data channels and other applications. Mathematics and Computers in Simulation, 19(3):183–191, October 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547790057X.

Richter:1977:NSL

[311] G. R. Richter. Numerical solution of Laplace’s equation as an integral equation of the first kind. Mathematics and Computers in Simulation, 19 (3):192–197, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900581.

Miller:1977:TDD

[312] A. J. Miller and P. Mars. Theory and design of a digital stochastic computer random number generator. Mathematics and Computers in Simulation, 19 (3):198–216, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900593.

Smith:1977:SCT

[313] M. W. A. Smith and A. P. Roberts. A study in continuous time of the identification of initial conditions and/or parameters of deterministic system by means of a Kalman-type filter. Mathematics and Computers in Simulation, 19(3):217–226, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847547790060X.

Anonymous:1977:Bb

[314] Anonymous. Bibliography. Mathematics and Computers in Simulation, 19 (3):227–237, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900611.

Anonymous:1977:E

[315] Anonymous. Erratum. Mathematics and Computers in Simulation, 19 (3):238, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900623.

Anonymous:1977:CPI

[316] Anonymous. Call for papers: IMACS-Symposium on simulation of control systems with special emphasis on modelling and redundancy. Mathematics and Computers in Simulation, 19(3):239, October 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900635.

Anonymous:1977:PO

[317] Anonymous. Pages 183–239 (October 1977). Mathematics and Computers in Simulation, 19(3):??, October 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wirz:1977:ICF

[318] H. J. Wirz, F. De Schutter, and A. Turi. An implicit, compact, finite difference method to solve hyperbolic equations. Mathematics and Computers in Simulation, 19(4): 241–261, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900428.

Arya:1977:MMC

[319] Suresh C. Arya, Roland P. Drewyer, and George Pincus. Mathematical modeling and computer simulation of elevated foundations supporting vibrating machinery. Mathematics and Computers in Simulation, 19 (4):262–271, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547790043X.

Sameh:1977:PDL

[320] Ahmed H. Sameh and David J.

Kuck. Parallel direct linear system solvers – a survey. Mathematics and Computers in Simulation, 19 (4):272–277, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900441.

Handler:1977:APC

[321] Wolfgang Händler. Aspects of parallelism in computer architecture. Mathematics and Computers in Simulation, 19(4):278–283, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900453.

Anonymous:1977:BRG

[322] Anonymous. Book review: Groundwater pollution: J. J. Fried, 1975, 338 pages, Elsevier Scientific Publishing Company, P.O. Box 211, Amsterdam, The Netherlands, Dfl. 95,00. Mathematics and Computers in Simulation, 19(4):284, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900465.

Anonymous:1977:BRM

[323] Anonymous. Book review: Multiobjective optimization in water resources systems: Y. Y. Haimes, W. A. Hall and H. T. Freedman, 1975, 200 pages, Elsevier Scientific Publishing Company, P.O. Box 211, Amsterdam, The Netherlands, Dfl. 42,00. Mathematics and Computers in Simulation, 19(4):284, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900477.

Anonymous:1977:BRB

[324] Anonymous. Book review: Basic linear partial differential equations: F. Treves, 1975, 496 pages, Academic Press Inc., 111 Fifth Avenue, New York, USA, US $29.50. Mathematics and Computers in Simulation, 19 (4):284–285, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900489.

Anonymous:1977:BRHa

[325] Anonymous. Book review: Handbook of circuit analysis languages and techniques: R. W. Jensen and L. P. Mc-Namee, 1976, 809 pages, Englewood Cliffs, NJ, Prentice Hall, US $34.50. Mathematics and Computers in Simulation, 19(4):285, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900519.

Anonymous:1977:BRHb

[326] Anonymous. Book review: The heat equation: D. V. Widder, 1975, 256 pages, Academic Press Inc., 111 Fifth Avenue, New York, USA, US $22.50. Mathematics and Computers in Simulation, 19(4):285, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900490.

Anonymous:1977:BRN

[327] Anonymous. Book review: Numerical solution of partial differential equations III: B. Hubbard (Editor), Proceedings of the symposium on the numerical solution of partial differential equations, synspade 1975, 1976, 509 pages, Academic Press Inc., US $25.00 or £13.00. Mathematics and Computers in Simulation, 19(4):285, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900507.

Anonymous:1977:BRP

[328] Anonymous. Book review: Progress in operations research: A. Prekopa (Editor), Colloquia Mathematica Societatis Janos Bolyai, 12., 1976, 968 pages in 2 volumes. North-Holland Publishing Company, US $73.25 or Dfl. 190.00. Mathematics and Computers in Simulation, 19(4):285, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900520.

Anonymous:1977:BRS

[329] Anonymous. Book review: Simulation of systems: L. Dekker (Editor), Proceedings of the 8th AICA Congress, Delft, The Netherlands, 23–28 August 1976. North Holland Publishing Company, Amsterdam, New York, 1977, 1122 pages, Dfl. 225.00, US $91.95. Mathematics and Computers in Simulation, 19(4):285–286, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900532.

VanRemoortere:1977:Bb

[330] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 19(4):287–295, December 1977. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475477900544.

Anonymous:1977:AIV

[331] Anonymous. Author index to volume XIX. Mathematics and Computers in Simulation, 19(4):296, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475477900556.

Anonymous:1977:SIV

[332] Anonymous. Subject index to volume XIX. Mathematics and Computers in Simulation, 19(4):297, December 1977. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475477900568.

Anonymous:1977:PD

[333] Anonymous. Pages 241–298 (December 1977). Mathematics and Computers in Simulation, 19(4):??, December 1977. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tomovic:1978:MCL

[334] Rajko Tomović. Modeling and control of large systems. Mathematics and Computers in Simulation, 20(1):1–4, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900472.

Ouchakov:1978:HCS

[335] V. B. Ouchakov and G. M. Petrov. Hybrid computing systems of the third generation. Mathematics and Computers in Simulation, 20(1):5–9, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900484.

Saphier:1978:DSM

[336] David Saphier. Dynamic simulation of a magnetohydrodynamic liquid-metal direct-current generator. Mathematics and Computers in Simulation, 20 (1):10–17, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900496.

Garfinkel:1978:CMR

[337] David Garfinkel, Michael C. Kohn, Murray J. Achs, John Phifer, and Gruia-Catalin Roman. Construction of more reliable complex metabolic models without repeated solution of their constituent differential equations. Mathematics and Computers in Simulation, 20(1):18–27, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900502.

Renner:1978:CST

[338] Gábor Renner. Computer simulation of the temperature and electromagnetic field in an inductively heated semiconductor. Mathematics and Computers in Simulation, 20 (1):28–36, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900514.

Alt:1978:EPF

[339] René Alt. Error propagation in Fourier transforms. Mathematics and Computers in Simulation, 20 (1):37–43, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900526.

Becker:1978:OSL

[340] Monique Becker and Pierre Becker. Optimal simulation lengths for various algorithms computing the mean. Mathematics and Computers in Simulation, 20(1):44–52, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900538.

Duket:1978:ESO

[341] Steven D. Duket and A. Alan B.

Pritsker. Examination of simulation output using spectral methods. Mathematics and Computers in Simulation, 20(1):53–60, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547890054X.

Gawler:1978:ESE

[342] R. A. Gawler and K. Morsztyn. Electronic simulation of electrical power system loads. Mathematics and Computers in Simulation, 20(1):61– 68, March 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900551.

Anonymous:1978:BRFa

[343] Anonymous. Book review: Foundations of microprogramming: architecture software, and application: A. K.

Agrawala and T. G. Rauscher: Academic Press, New York, 1976, 416 pp., $15.50. Mathematics and Computers in Simulation, 20(1):69, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900575.

Anonymous:1978:BRIa

[344] Anonymous. Book review: Introduction to microcomputers and microprocessors: A. Barna and D. Porat: Wiley, New York, 1976, 108 pp., $10.50. Mathematics and Computers in Simulation, 20(1):69, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900587.

Anonymous:1978:BRMa

[345] Anonymous. Book review: Modelling and performances evaluation of computer systems: H. Beilner and E. Gelenbe (Editors): North Holland Publishing Company, (NL or USA), 514 pages, US$38.95 or Dfl. 95.00, 1977. Mathematics and Computers in Simulation, 20(1):69, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900563.

Anonymous:1978:BRDa

[346] Anonymous. Book review: Digital circuits and logic design: S. C. Lee: 1976, 594 pages Englewood Cliffs, NJ, Prentice Hall, $24.00. Mathematics and Computers in Simulation, 20(1):69–70, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900599.

Anonymous:1978:BRDb

[347] Anonymous. Book review: D. F.

Stout and M. Kaufman: Handbook of operational amplifier circuit design: 1976, New York, McGraw-Hill, $24.50. Mathematics and Computers in Simulation, 20(1):70, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900617.

Anonymous:1978:BREa

[348] Anonymous. Book review: Engineering simulation using small scientific computers: M. J. Shah: (Automatic computation series), Prentice Hall, Englewood Cliffs, NJ., 1976, 401 pp., $14.95. Mathematics and Computers in Simulation, 20(1):70, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900605.

Anonymous:1978:Ba

[349] Anonymous. Bibliography. Mathematics and Computers in Simulation, 20(1):71–80, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900629.

Anonymous:1978:ARG

[350] Anonymous. Article review: G. Ram, M.I.S. — The METEOSTAT Image Simulator, Mathematics and Computers in Simulation XIX (1977) 151–158.

Mathematics and Computers in Simulation, 20(1):81, March 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900630.

Anonymous:1978:EB

[351] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 20(1):ifc, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900460.

Anonymous:1978:PM

[352] Anonymous. Pages 1–82 (March 1978). Mathematics and Computers in Simulation, 20(1):??, March 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Chang:1978:OCF

[353] P. W. Chang and B. A. Finlayson. Orthogonal collocation on finite elements for elliptic equations. Mathematics and Computers in Simulation, 20(2):83–92, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900319.

Miranker:1978:PMS

[354] W. L. Miranker. Parallel methods for solving equations. Mathematics and Computers in Simulation, 20(2):93–101, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900320.

Smith:1978:EEB

[355] M. W. A. Smith and A. P. Roberts.

An exact equivalence between the discrete- and continuous-time formulations of the Kalman filter. Mathematics and Computers in Simulation, 20(2):102–109, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900332.

Kaplan:1978:SGT

[356] B. Z. Kaplan and Y. Tatrash. Stabilized generation of three phase sinewaves. Mathematics and Computers in Simulation, 20(2):110–113, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900344.

Eltimsahy:1978:CSS

[357] Adel H. Eltimsahy and Clifford H.

Copass. Computer simulation for a solar heating system. Mathematics and Computers in Simulation, 20 (2):114–127, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900356.

Anonymous:1978:BRCa

[358] Anonymous. Book review: Computer Networking: R. P. Blanc and I. W. Cotton (Ed.), IEEE Service Center, 445 Hoes Lane, Piscataway, N.J. 08854, U.S.A., 1976, US $17.95. Mathematics and Computers in Simulation, 20(2):128, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900368.

Anonymous:1978:BRFb

[359] Anonymous. Book review: Filtering in the time and frequency domains: H. J. Blinchikoff and A. I. Zuerev, 1976, 494 pages, Wiley, New York, $27.50. Mathematics and Computers in Simulation, 20(2):128, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547890037X.

Anonymous:1978:BRMb

[360] Anonymous. Book review: Multivariable computer control: Fisher, D. G. and Seborg, D. E., A case study. North Holland Publishing Company, P.O. Box 211, Amsterdam, The Netherlands, 1976, 205 pages, Dfl. 60.00. Mathematics and Computers in Simulation, 20(2):128, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900381.

Anonymous:1978:BRCb

[361] Anonymous. Book review: Computer simulation of water resources systems: Vansteenkiste G. (Editor), Proceedings of the First IFIP Working conference on computer simulation of water resource systems, Ghent, Belgium, 30 July–2 August 1974, North Holland Publishing Company, 1975, 680 pages, Dfl. 135.000. Mathematics and Computers in Simulation, 20(2):129, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900423.

Anonymous:1978:BREb

[362] Anonymous. Book review: Economic modeling for water policy evaluation: Thrall R. M., Thomson R., Schad T. and Heady E., North Holland Publishing Company, 1976, 262 pages, Dfl. 54.00. Mathematics and Computers in Simulation, 20(2):129, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900411.

Anonymous:1978:BRMc

[363] Anonymous. Book review: Microprocessors: Technology, architecture and applications: D. R. McGlynn, 1976, 207 pages, Wiley, New York, $11.95. Mathematics and Computers in Simulation, 20(2):129, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547890040X.

Anonymous:1978:BRU

[364] Anonymous. Book review: The User’s Handbook of D/A and A/D Converters: E. R. Hnatek, 1976, 472 pages, Wiley, New York, $24.95. Mathematics and Computers in Simulation, 20(2):129, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900393.

Anonymous:1978:BRSa

[365] Anonymous. Book review: System stimulation in water resources:

Vansteenkiste G. (Editor), Proceedings of the IFIP Working Conference held at Bruges, Belgium from September 3–5, 1975. North Holland Publishing Company, 1976, Dfl. 105.000. Mathematics and Computers in Simulation, 20(2):129–130, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900435.

Anonymous:1978:Bb

[366] Anonymous. Bibliography. Mathematics and Computers in Simulation, 20(2):131–142, June 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900447.

Anonymous:1978:BRAa

[367] Anonymous. Book review: Aspects of parallelism in computer architecture: W. Händler, Mathematics and Computers in Simulation XIX (1977) 278–283. Mathematics and Computers in Simulation, 20(2):143–145, June 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900459.

Anonymous:1978:PJ

[368] Anonymous. Pages 83–146 (June 1978). Mathematics and Computers in Simulation, 20(2):??, June 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Saaty:1978:MUD

[369] Thomas L. Saaty. Modeling unstructured decision problems — the theory of analytical hierarchies. Mathematics and Computers in Simulation, 20(3):147–158, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900642.

Simon:1978:UMS

[370] Herbert A. Simon. The uses of mathematics in the social sciences. Mathematics and Computers in Simulation, 20(3):159–166, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900654.

Huyakorn:1978:PEF

[371] Pongsarl Huyakorn and George F.

Pinder. A pressure-enthalpy finite element model for simulating hydrothermal reservoirs. Mathematics and Computers in Simulation, 20(3): 167–178, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900666.

Zegwaard:1978:HSN

[372] J. F. Zegwaard and W. Veldhuyzen.

Hybrid simulation of a nonlinear helmsman model. Mathematics and Computers in Simulation, 20(3):179– 189, September 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900678.

Carver:1978:EID

[373] M. B. Carver. Efficient integration over discontinuities in ordinary differential equation simulations. Mathematics and Computers in Simulation, 20(3):190–196, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847547890068X.

Csendes:1978:NFE

[374] Z. J. Csendes. A novel finite element method for two-point boundary value problems. Mathematics and Computers in Simulation, 20(3): 197–203, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900691.

Shampine:1978:SOD

[375] L. F. Shampine. Solving ordinary differential equations for simulation. Mathematics and Computers in Simulation, 20(3):204–207, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900708.

Nevanlinna:1978:NSM

[376] O. Nevanlinna. A note on semidiscretization methods. Mathematics and Computers in Simulation, 20(3): 208–210, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847547890071X.

Anonymous:1978:BRCc

[377] Anonymous. Book review: Computeraided experimentation: Interfacing to minicomputers: J. Finkel 1975, 422 pages, Wiley, New York. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900770.

Anonymous:1978:BRCd

[378] Anonymous. Book review: Computer interfacing and on-line operation: J. C. Cluley 1975, 181 pages, Crane Russak, New York. Mathematics and Computers in Simulation, 20(3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900745.

Anonymous:1978:BRG

[379] Anonymous. Book review: Global simulation models: a comparative study: J. Clark and S. Coles 1975, 135 pages, Wiley, London. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900733.

Anonymous:1978:BRIb

[380] Anonymous. Book review: Introduction to digital filtering: R. E. Bogner and A. G. Constantinides 1975, 198 pages, Wiley, London. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900721.

Anonymous:1978:BRMd

[381] Anonymous. Book review: Modeling and control of kraft production systems: I. J. Williams 1975, 284 pages, I.S.A., Pittsburg. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900800.

Anonymous:1978:BRMe

[382] Anonymous. Book review: Models for flow systems and chemical reactors: C. Y. Wen and L. T. Fan 1975, Marcel Dekker, New York, 570 pages. Mathematics and Computers in Simulation, 20(3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900794.

Anonymous:1978:BRPa

[383] Anonymous. Book review: Proceedings of the 7th IFIP conference on optimization techniques: Modeling and optimization in the service of man. Part I and II: J. CEA (Ed.) 1976, Springer-Verlag, Berlin, 1706 pages. Mathematics and Computers in Simulation, 20(3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900824.

Anonymous:1978:BRSb

[384] Anonymous. Book review: A system approach to water disturbation modeling and control: R. Demoyer and L. B. Horwitz 1975, 143 pages, Lexington Books, Lexington. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900757.

Anonymous:1978:BRSc

[385] Anonymous. Book review: Solution of boundary value problems by the method of integral operators: D. L. Colton 1976, 160 pages, $10.00. Fearon-Pitman Publishers Inc., 6 Davis Drive, Belmont, CA 94002 USA. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900836.

Anonymous:1978:BRSd

[386] Anonymous. Book review: Simulation with GASP PL/1: A. B. Pritsker and R. E. Young 1975, 335 pages, Wiley, New York. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900782.

Anonymous:1978:BRSe

[387] Anonymous. Book review: Sparsity — its practical application to systems analysis: A. Brameller, R. N. Allan and Y. M. Hamam 1976, 180 pages, $15.90. Fearon-Pitman Publishers Inc. 6 Davis Drive, Belmont, CA 94002 USA. Mathematics and Computers in Simulation, 20(3):211, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900812.

Anonymous:1978:BRV

[388] Anonymous. Book review: A vector space approach to models and optimization: C. N. Dorny 1975, 599 pages, Wiley, New York. Mathematics and Computers in Simulation, 20 (3):211, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900769.

Anonymous:1978:BREc

[389] Anonymous. Book review: Elements of system theory: P. Faurre and M. Depeyrot 1976, 282 pages, North Holland Publishing Company, Amsterdam and New York, US $24.95 — Dfl. 65.00. Mathematics and Computers in Simulation, 20(3):211–212, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900848.

Anonymous:1978:BRCe

[390] Anonymous. Book review: Computerized operation of power systems: S. C. Savulescu 1976, 302 pages, Elsevier, Amsterdam. Mathematics and Computers in Simulation, 20 (3):212, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900885.

Anonymous:1978:BRDc

[391] Anonymous. Book review: Dynamic stochastic models from empirical data: R. L. Kashyap and A. R. Rao 1976, 334 pages, Academic Press, London. Mathematics and Computers in Simulation, 20(3):212, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900873.

Anonymous:1978:BRFc

[392] Anonymous. Book review: Function theoretic methods in differential equations: R. Gilbert and R. Wilnatch (Editors) 1976, 320 pages, $17.00. Fearon-Pitman Publishers Inc., 6 Davis Drive, Belmot, CA 94002 USA. Mathematics and Computers in Simulation, 20 (3):212, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547890085X.

Anonymous:1978:BRIc

[393] Anonymous. Book review: Introduction to operations research: a computer oriented algorithmic approach: B. E. Gillet 1976, 617 pages, Mac Graw Hill, New York. Mathematics and Computers in Simulation, 20 (3):212, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900861.

Anonymous:1978:BRPb

[394] Anonymous. Book review: Parallel computers — parallel mathematics: M. Feilmeier (Editor) Proceedings of the IMACS — GI symposium of March 14–16, 1977, in Munich. 354 pages, Dfl. 90.00. North Holland Publishing Company, Amsterdam, New York, Oxford. Mathematics and Computers in Simulation, 20 (3):212–213, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900897.

Anonymous:1978:BRAb

[395] Anonymous. Book review: Algorithmic methods in probability: M. F. Neuts (Edt.) 1977, 296 pages, Dfl. 65.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 20(3):213, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475478900903.

Anonymous:1978:BRAc

[396] Anonymous. Book review: Advances in operations research: M. Roubens (Editor) Proceedings of the EURO II; the second European congress on Operations research, Stockholm, 1976, November 29, December 1, 667 pages, 1977, Dfl. 145.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 20(3):213, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900915.

Anonymous:1978:Bc

[397] Anonymous. Bibliography. Mathematics and Computers in Simulation, 20(3):214–226, September 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900927.

Anonymous:1978:PS

[398] Anonymous. Pages 147–226 (September 1978). Mathematics and Computers in Simulation, 20(3):??, September 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vignes:1978:NME

[399] J. Vignes. New methods for evaluating the validity of the results of mathematical computations. Mathematics and Computers in Simulation, 20(4):227–249, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900162.

Amouroux:1978:HSA

[400] M. Amouroux, J. P. Babary, A. El Jai, and J. P. Gouyon. Hybrid simulation of adaptive open loop control for parabolic systems. Mathematics and Computers in Simulation, 20(4): 250–258, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900174.

Saidel:1978:SDM

[401] Gerald M. Saidel, Paramjit S. Chandhoke, and Mark A. Knepper. Spatially discrete models of countercurrent mass transport for application to the kidney. Mathematics and Computers in Simulation, 20(4): 259–270, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900186.

Tu:1978:CDF

[402] Yih-O Tu. Calculation of demagnetizing field by means of FFT. Mathematics and Computers in Simulation, 20(4):271–284, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900198.

Kobayashi:1978:FPA

[403] Yasuhiro Kobayashi, Masaaki Ohkita, and Michio Inoue. Fractional power approximations of elliptic integrals and Bessel functions. Mathematics and Computers in Simulation, 20(4): 285–290, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900204.

Smith:1978:KFT

[404] M. W. A. Smith and A. P. Roberts. A Kalman filter type of extension to a deterministic gradient technique for parameter estimation. Mathematics and Computers in Simulation, 20(4): 291–300, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900216.

Hasse:1978:IMS

[405] T. Hasse. An iterative method for solving an exact initial value problem of the second order by a hybrid computer. Mathematics and Computers in Simulation, 20(4):301– 307, December 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900228.

Auslander:1978:CSS

[406] David M. Auslander. A continuoussystem simulation language designed for LSI economics. Mathematics and Computers in Simulation, 20(4): 308–313, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547890023X.

Anonymous:1978:BRCf

[407] Anonymous. Book review: Computer architecture: G. G. Boulaye and D. W. Lewin: Transactions of the NATO Advanced Study Institute, held in St. Raphael, France, 12–24 September 1976. Published by D. Reidel Publishing Company, Dordrecht-Holland/ Boston-USA. in 1977. Dfl. 70.- or US $28.00. ISBN 90-277-0803-7. Mathematics and Computers in Simulation, 20(4):314, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900241.

Anonymous:1978:BRId

[408] Anonymous. Book review: Information theory, statistical decision functions, random processes: J. Kozesnik: Transaction of the Seventh Prague Conference on Information Theory, Statistical Decision Functions, Random Processes and of the 1974 European meeting of Statisticians, held at Prague, August 18– 23, 1974. Published by D. Reidel Publishing Company, Dordrecht-Holland or Boston-USA. Dfl/130.00 or US $52.00 for Volume A; ISBN 90-277-0852-5. Mathematics and Computers in Simulation, 20(4):314, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900253.

Anonymous:1978:BRSf

[409] Anonymous. Book review: Structure of Interconnected Systems: H. Nicholson: IEE Control Engineering Series 5. Published in March 1978 by Peter Peregrinus Ltd. on behalf of the Institution of Electrical Engineers (South Gate House, Stevenage, Herts SG1 1HQ England). $12.00 — ISBN 0-901223-69-7. Mathematics and Computers in Simulation, 20(4):314, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900265.

Anonymous:1978:BRPc

[410] Anonymous. Book review: Proceedings of the 1978 UKSC Conference on computer simulation. 4–6 April 1978, Grosverhor Hotel, Chester, England: UKSC 78: Published by IPC Science and Technology Press, IPC House, 32 High Street, Guilford, Surrey, England GU1 3EW. ISBN 0-90285-92-2. Mathematics and Computers in Simulation, 20(4):314–315, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900277.

Anonymous:1978:Bd

[411] Anonymous. Bibliography. Mathematics and Computers in Simulation, 20(4):316–328, December 1978. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900289.

Anonymous:1978:SIV

[412] Anonymous. Subject index to volume XX. Mathematics and Computers in Simulation, 20(4):329, December 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900290.

Anonymous:1978:AIV

[413] Anonymous. Author index to volume XX. Mathematics and Computers in Simulation, 20(4):330–331, December 1978. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475478900307.

Anonymous:1978:PD

[414] Anonymous. Pages 227–331 (December 1978). Mathematics and Computers in Simulation, 20(4):??, December 1978. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Saaty:1979:AAH

[415] Thomas L. Saaty. Applications of analytical hierarchies. Mathematics and Computers in Simulation, 21(1):1–20, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901010.

Mars:1979:HSS

[416] P. Mars, F. G. Mcintosh, and T. Baxter. High-speed simulation of discrete dynamic probabilistic systems. Mathematics and Computers in Simulation, 21(1):21–38, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901022.

Wong:1979:NOI

[417] J. K. Wong. A note on the optimal inverter problem. Mathematics and Computers in Simulation, 21 (1):39–41, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901034.

Badel:1979:GNA

[418] Marc Badel. Gènèration de nombres alèatoires corrèlès. (French) [Generation of correlated random numbers].

Mathematics and Computers in Simulation, 21(1):42–49, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901046.

Leroudier:1979:DES

[419] Jacques Leroudier and Michel Parent. Discrete event simulation modelling of computer systems for performance evaluation. Mathematics and Computers in Simulation, 21 (1):50–79, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901058.

Metz:1979:NMA

[420] Michel Metz, Henri Foch, and Bernard Trannoy. New method of analysing static converter behaviour using global analogue simulation. Mathematics and Computers in Simulation, 21 (1):80–84, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990106X.

vandeVen:1979:SPE

[421] H. H. van de Ven and A. W. M.

Groenendaal. System parameters estimation by automatic adjustment of an analogue model. Mathematics and Computers in Simulation, 21 (1):85–93, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901071.

Maurin:1979:SAC

[422] T. Maurin and C. Sol. Simulation analogique d’un convertisseur électrique statique à l’aide d’un calculateur réversible. (French) [analog simulation of a static electric converter using a reversible calculator]. Mathematics and Computers in Simulation, 21(1):94–101, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901083.

Khalil:1979:CGD

[423] Hatem M. Khalil. Computer generation of difference approximations. Mathematics and Computers in Simulation, 21(1):102–108, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901095.

Eitelberg:1979:NSS

[424] Eduard Eitelberg. Numerical simulation of stiff systems with a diagonal splitting method. Mathematics and Computers in Simulation, 21 (1):109–115, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901101.

Anonymous:1979:BRCa

[425] Anonymous. Book review: Computation of power system transients: J. P. Bickford et al. 1976, International Scholarity Book Services, Forest Grove, OR., USA, £27.50. 176 pages. Mathematics and Computers in Simulation, 21(1):116, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901125.

Anonymous:1979:BRI

[426] Anonymous. Book review: Improperly posed boundary value problems: A. Carasso and A. P. Stone (Editors): 1975, 160 pages, £10.00. Fearon-Pitman Publishers Inc., 6 Davis Drive, Belmont, CA 94002 USA. Mathematics and Computers in Simulation, 21(1):116, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901113.

Anonymous:1979:BREa

[427] Anonymous. Book review: Elements of system theory: P. Faurre and M. Depeyrot: 1976, 282 pages, Dfl. 65.00, North Holland Publishing Company, Amsterdam, New York. Mathematics and Computers in Simulation, 21 (1):116–117, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901137.

Anonymous:1979:BRAa

[428] Anonymous. Book review: Applied computation theory: Analysis, design, modeling: R. T. Yeh: 1976, 624 pages, Prentice-Hall, Engelwood Cliffs, N.J., USA, £22.50. Mathematics and Computers in Simulation, 21(1):117, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901162.

Anonymous:1979:BRCb

[429] Anonymous. Book review: The complexity of computing: J. E. Savage: 391 pages, 1976, J. Wiley, New York, £22.95, ISBN 0-471-75517-6. Mathematics and Computers in Simulation, 21(1):117, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901150.

Anonymous:1979:BRO

[430] Anonymous. Book review: Operational research 75: K. B. Haley (Editor): Proceedings of the Seventh IFORS International Conference on Operational Research, Tokyo, July 17–23, 1975. 698 pages, 1976, Dfl. 140.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 21(1):117, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901149.

Anonymous:1979:BRT

[431] Anonymous. Book review: Theory of modelling and simulation: B. P. Zeigler: 435 pages, J. Wiley, New York, £24.95, ISBN 0-471-98152-4. Mathematics and Computers in Simulation, 21(1):117–118, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901174.

Anonymous:1979:BRMa

[432] Anonymous. Book review: Mathematical programming methods: G. Zoutendijk: 300 pages, 1976, American Elsevier, New York, £27.50, ISBN 0-444-11069-0. Mathematics and Computers in Simulation, 21(1):118, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901186.

Anonymous:1979:BRMb

[433] Anonymous. Book review: Manpower planning models: R. C. Grinold and K. T. Marshall: Modelling of long range manpower in large organizations. 1977. 256 pages, Dfl. 65.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 21(1):118, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901198.

Anonymous:1979:BRF

[434] Anonymous. Book review: Fuzzy automata and decision processes: M. M. Gupta and G. N. Saridis (Editors): 1977, 336 pages, Dfl. 75.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 21 (1):118–119, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901204.

Anonymous:1979:BRAb

[435] Anonymous. Book review: Advances in computer methods for partial differential equations II: R. Vichnevetsky(Editor): IMACS Symposium, The Rutgers university, Department of Computer Science, New Brunswick, New Jersey, 08903, U.S.A., 389 pages, 1977. Mathematics and Computers in Simulation, 21(1):119, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990123X.

Anonymous:1979:BRSa

[436] Anonymous. Book review: Stochastic integration and generalized martingales: A. U. Kussmaul: 1977, 208 pages, £12.75. Fearon-Pitman Publishers Inc., 6 Davis Drive, Belmont, CA 94002 USA. Mathematics and Computers in Simulation, 21(1):119, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901228.

Anonymous:1979:BRSb

[437] Anonymous. Book review: Social systems design, normative theory and the MAPS design technology: R. H. Kilmann: 1977, 328 pages, Dfl. 65.00. North Holland Publishing Company, Amsterdam and New York. Mathematics and Computers in Simulation, 21(1):119, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901216.

Anonymous:1979:BRN

[438] Anonymous. Book review: Numerical methods for differential equations and simulation: A. W. Bennett and R. Vichnevetsky (Editors): North Holland Publishing Company, The Netherlands 220 Pages, IBSN 0-444-85111-9, 1978. Mathematics and Computers in Simulation, 21(1):119–120, March 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901241.

Anonymous:1979:Ba

[439] Anonymous. Bibliography. Mathematics and Computers in Simulation, 21(1):121–136, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901253.

Anonymous:1979:EB

[440] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 21(1):ifc, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901009.

Anonymous:1979:PM

[441] Anonymous. Pages 1–136 (March 1979).

Mathematics and Computers in Simulation, 21(1):??, March 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Beljakov:1979:MEE

[442] V. G. Beljakov. Methods of evaluating the equivalent speed of analog processors. Mathematics and Computers in Simulation, 21(2):137– 140, August 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901265.

Mohapatra:1979:MCT

[443] Pratap K. J. Mohapatra. Modal control theory and industrial dynamics — a dual approach to the design of policy decisions in a multistage production-inventory system. Mathematics and Computers in Simulation, 21(2):141–149, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901277.

Hoeld:1979:NLD

[444] A. Hoeld and D. Beraha. A nonlinear digital simulator for BWR nuclear power plants and its application within the scope of optimal digital load-following control. Mathematics and Computers in Simulation, 21 (2):150–162, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901289.

Ur:1979:FEI

[445] Hanoch Ur and Chaim D. Shen-Orr. Function evaluation by incremental computation, with applications. Mathematics and Computers in Simulation, 21(2):163–169, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901290.

Vichnevetsky:1979:SCN

[446] R. Vichnevetsky. Stability charts in the numerical approximation of partial differential equations: a review. Mathematics and Computers in Simulation, 21(2):170–177, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901307.

Simionescu:1979:CHS

[447] Claudia Lidia Simionescu. Control for an hereditary system. Mathematics and Computers in Simulation, 21 (2):178–181, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901319.

Neuschl:1979:PDM

[448] S̆. Neuschl and I. Burger. Parallel digital models of continuous systems and their application to digital filters synthesis. Mathematics and Computers in Simulation, 21(2):182–188, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901320.

Lipitakis:1979:NFP

[449] E. A. Lipitakis and D. J. Evans.

Normalized factorization procedures for the solution of self-adjoint Elliptic partial differential equations in three-space dimensions. Mathematics and Computers in Simulation, 21 (2):189–196, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901332.

Meisinger:1979:SMV

[450] Reinhold Meisinger. Simulation of maglev vehicles riding over single and double span guideways. Mathematics and Computers in Simulation, 21 (2):197–206, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901344.

Kaplan:1979:NSG

[451] B. Z. Kaplan and S. T. Bachar. A new stabilized generator model for three phase sinewaves. Mathematics and Computers in Simulation, 21 (2):207–208, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901356.

Hopmans:1979:ISS

[452] A. C. M. Hopmans and J. P. C. Kleijnen. Importance sampling in systems simulation: a practical failure? Mathematics and Computers in Simulation, 21(2):209–220, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901368.

Parthasarathy:1979:LMR

[453] S. Parthasarathy and P. Deschizeaux. A logical model for the remote control of power system topology. Mathematics and Computers in Simulation, 21(2):221–225, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990137X.

Kobayashi:1979:UEF

[454] Y. Kobayashi, M. Ohkita, and M. Inoue.

On the use of an exponential function in approximation of elliptic integrals. Mathematics and Computers in Simulation, 21(2):226–230, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901381.

Vergnes:1979:FOP

[455] J. Vergnes and J. Dumontet. Finding an optimal partition for a numerical integration using the trapezoidal rule. Mathematics and Computers in Simulation, 21(2):231–241, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901393.

Anonymous:1979:BRAc

[456] Anonymous. Book review: Annotated bibliographies of simulation: Simulation Councils Proceedings Series. Volume 4, Number 1 — June 1974 (S.C.S.

— P.O. Box 2228 — La Jolla California 92038 USA). 132 pages. Mathematics and Computers in Simulation, 21(2):242, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990140X.

Anonymous:1979:BRSc

[457] Anonymous. Book review: Stochastic systems — vol. 1: Modeling, identification; vol. 2: Optimization: Mathematical Programming Studies Vol. 5 and 6. North-Holland Publishing Company, The Netherlands, 1976, Vol. 1: 244 pages $21.25. Vol. 2: 264 pages $21.25. Mathematics and Computers in Simulation, 21(2):242, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901411.

Anonymous:1979:BRPa

[458] Anonymous. Book review: Power system control and stability: Iowa State University Press, Ames, IA, USA, 1977, pages, $45.00. Mathematics and Computers in Simulation, 21 (2):242–243, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901423.

Anonymous:1979:BRCc

[459] Anonymous. Book review: Computer science and multiple-values logic: Theory and applications: North-Holland Publishing Company, The Netherlands, 1977, 548 pages, $55.50. ISBN 0-7204-0406-1. Mathematics and Computers in Simulation, 21(2): 243, August 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901459.

Anonymous:1979:BREb

[460] Anonymous. Book review: Evaluation of policy simulation models: a conceptual approach and case study: Information Resources Press, Washington, 1977, 350 pages, $30.50. ISBN 0-87815-018-8. Mathematics and Computers in Simulation, 21(2):243, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901447.

Anonymous:1979:BRPb

[461] Anonymous. Book review: Parameter estimation in engineering and science: New York, Wiley, 1977, 501 pages. $24.95, ISBN 0-471-06118-2. Mathematics and Computers in Simulation, 21(2):243, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901435.

Anonymous:1979:BRMc

[462] Anonymous. Book review: Mathematical modeling and digital simulation for engineers and scientists: New York, Wiley, 1977, 332 pages, $21.00. ISBN 0-471-80344-8. Mathematics and Computers in Simulation, 21(2):243–244, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479901460.

Anonymous:1979:BRMd

[463] Anonymous. Book review: Modern network theory: an introduction: Georgi Publishing Company, St-Saphorin, Switzerland, 1978, 267 pages, SF 63.00. ISBN 2-604-00034-2. Mathematics and Computers in Simulation, 21(2):244, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901472.

Anonymous:1979:BRD

[464] Anonymous. Book review: Discrete and switching functions: Georgi Publishing Company, St-Saphorin, Switzerland, and McGraw-Hill International Book Company, 1978, 729 pages, ISBN 0-07-015509-7. Mathematics and Computers in Simulation, 21(2): 244–245, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901484.

Anonymous:1979:BRCd

[465] Anonymous. Book review: Circuit theory and design: Georgi Publishing Company, St. Saphorin, Switzerland, 1978, 624 pages, SF 95.00. ISBN 2-604-00033-4. Mathematics and Computers in Simulation, 21(2):245, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901496.

VanRemoortere:1979:B

[466] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 21(2):246–261, August 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479901502.

Anonymous:1979:PA

[467] Anonymous. Pages 137–261 (August 1979). Mathematics and Computers in Simulation, 21(2):??, August 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zboril:1979:SPP

[468] František Zbořil. Solving parabolic PDE’s by a decomposition method. Mathematics and Computers in Simulation, 21(3):263–269, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475479900715.

Evans:1979:DMS

[469] D. J. Evans. Direct methods of solution of partial differential equations with periodic boundary conditions. Mathematics and Computers in Simulation, 21(3):270–275, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900727.

Korn:1979:PCS

[470] Granino A. Korn and Gunther Lafleur.

Programming continuous-system simulation in PASCAL. Mathematics and Computers in Simulation, 21(3): 276–281, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900739.

Benson:1979:IMS

[471] A. Benson and D. J. Evans. Iterative methods of solution for the elliptic fourth boundary value problem. Mathematics and Computers in Simulation, 21(3):282–288, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900740.

Allison:1979:SHS

[472] J. S. Allison and H. M. Johnson. Stability of hybrid simulation of dynamic systems. Mathematics and Computers in Simulation, 21(3): 289–303, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900752.

VanRemoortere:1979:BRAa

[473] P. Van Remoortere. Book review: Automata, languages, development. At the crossroads of biology, mathematics and computer science: A. Lindenmayer and G. Rozenberg: North-Holland Publishing Company, Amsterdam and New York, 1976. ISBN 0-7204-0474-6. 529 pages. US $58.50. Mathematics and Computers in Simulation, 21(3):304, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900831.

VanRemoortere:1979:BRCa

[474] P. Van Remoortere. Book review: Computer simulation of water resources systems proceedings of the first IFIP working conference on modelling and simulation of water resources, Ghent, Belgium; July 30–August 2, 1974: G. C. Vansteenkiste: North-Holland Publishing Company, Amsterdam, 1975. ISBN 0-7204-2829-7. 686 pages, US $65.75. Mathematics and Computers in Simulation, 21(3):304, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900806.

VanRemoortere:1979:BRDa

[475] P. Van Remoortere. Book review: Developmental systems and languages: G. T. Herman and G. Rozenberg: North-Holland Publishing Company, Amsterdam and New York, 1975. ISBN 0-7204-2806-8. 366 pages. US $36.50. Mathematics and Computers in Simulation, 21(3):304, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900776.

VanRemoortere:1979:BRDb

[476] P. Van Remoortere. Book review:

Dynamic stochastic models from empirical data: R. L. Kashyap and A. R. Rao: Academic Press, New York and London. 1976. ISBN 0-12-400550-0. 352 pages, US $34.50. Mathematics and Computers in Simulation, 21(3):304, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990082X.

VanRemoortere:1979:BRIa

[477] P. Van Remoortere. Book review: Iterative structural design: W. R. Spillers: North-Holland Publishing Company, Amsterdam and New York, 1975. ISBN 0-7204-0368-5. 187 pages, US $21.25. Mathematics and Computers in Simulation, 21(3):304, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847547990079X.

VanRemoortere:1979:BRMa

[478] P. Van Remoortere. Book review: Man and computer: M. Marios (Editor): Proceedings of the second international conference, Bordeaux 11– 16 September 1972. North-Holland Publishing Company, Amsterdam and New York, 1975. ISBN 0-7204-0305-7. 610 pages. US $80.00. Mathematics and Computers in Simulation, 21 (3):304, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900788.

VanRemoortere:1979:BRMb

[479] P. Van Remoortere. Book review: Minicomputer software Proceedings of the IFIP conference on software for minicomputers, Lake Balaton, Hungary, September 8–12, 1975: J. R. Bell and C. G. Bell (editors): North-Holland Publishing Company, Amsterdam, 1976; ISBN 0-7204-0437-1. 334 pages, US $41.50. Mathematics and Computers in Simulation, 21 (3):304, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900818.

VanRemoortere:1979:BRPa

[480] P. Van Remoortere. Book review: Principles of continuous system simulation.

— analog, digital and hybrid simulation in a computer science perspective: Gilio: Teubner Studienbücher, ISBN 3-519-02336-9. 1975. 172 pages. DM 28.50.

Mathematics and Computers in Simulation, 21(3):304, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900764.

VanRemoortere:1979:BREa

[481] P. Van Remoortere. Book review: Economic modelling for water policy evaluation: R. M. Thrall, R. Thompson, T. Schad and E. Heady: North-Holland Publishing Company, Amsterdam and New York. 270 pages, 1976. ISBN 0-7204-0387-1. US $29.25. Mathematics and Computers in Simulation, 21 (3):304–305, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900843.

VanRemoortere:1979:BRT

[482] P. Van Remoortere. Book review: Theory of modelling and simulation: B. P. Ziegler: John Wiley and Sons, New-York, London, Sydney, Toronto, ISBN 0-471-98152-4. 435 pages, 1976. Mathematics and Computers in Simulation, 21(3):305–306, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900855.

VanRemoortere:1979:BRCb

[483] P. Van Remoortere. Book review:

Computer applications in medical care: IEEE Publications: Proceedings of the 1st symposium on computer application in medical care, Washington, October 3–5, 1977. 373 pages, US $ 25.00 — 77CH1270-8C. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900879.

VanRemoortere:1979:BRCc

[484] P. Van Remoortere. Book review: Computer networking: IEEE Publications: Proceedings of the computer networking symposium held in Gaithersburg, MD, in December 15, 1977. 130 pages. US$12.00—77CH1252-6C. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900922.

VanRemoortere:1979:BRCd

[485] P. Van Remoortere. Book review: Computer software and applications: IEEE Publications: Proceedings of the conference on computer software and applications, held November 8– 11, 1977 in Chicago, IL. 834 pages, US $25.00 — 77CH1291-4C. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900909.

VanRemoortere:1979:BRCe

[486] P. Van Remoortere. Book review: Computer science: IEEE Publications: Conference record of the 18th symposium of foundations of computer science, held October 31–November 2, 1977 in Providence, RI. 269 pages, US $20.00 — 77CH1278-1C. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900892.

VanRemoortere:1979:BRDc

[487] P. Van Remoortere. Book review:

Digital signal computers and processors: A. C. Salazar: a book of selected reprints, IEEE Press, 1977, 345 pages, US $25.95. ISBN 0-87942-099-5. Mathematics and Computers in Simulation, 21(3):306, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900934.

VanRemoortere:1979:BRMc

[488] P. Van Remoortere. Book review: Microprogramming (Micro 10).: IEEE Publications: Proceedings of the 10th workshop on microprogramming, held October 5–7, 1977 in Niagara Falls, NY. 133 pages, US $12.00 — 77CH1266-6. Mathematics and Computers in Simulation, 21(3):306, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900880.

VanRemoortere:1979:BRPb

[489] P. Van Remoortere. Book review: Parallel processing: IEEE Publications: Proceedings of the international conference on parallel processing held August 23– 26, 1977 in Bellaire, MI. 246 pages, US $20.00 — 77CH1253-4C. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900867.

VanRemoortere:1979:BRSa

[490] P. Van Remoortere. Book review: Simulation: IEEE Publications: Proceedings of the 9th winter conference on simulation held December 5–7, 1977 in Gaithersburg, MD. 880 pages, US $ 40.00 — 77CH1309-4SMC. Mathematics and Computers in Simulation, 21(3):306, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900910.

VanRemoortere:1979:BRMd

[491] P. Van Remoortere. Book review: Markov modells mit indirekter beobachtung: Dr. F. Benito: CEin algorithmus zur optimierung der steuerungs und beobachtungspolitik. — Schriften des Instituts für Operations Research des ETH Zürich 2. 1978. Publ. by Haupt, Falkenplatz 14, CH 3001, Bern. 230 pages. DM $38. — ISBN 3-258-02811-7. Mathematics and Computers in Simulation, 21(3): 306–307, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900946.

VanRemoortere:1979:BRAb

[492] P. Van Remoortere. Book review: Analog systems for microprocessors and minicomputers: P. H. Garrett: Reston Publishing Company, Inc., Reston Virginia 22090, USA, 272 pages, March 1978, $19.95. Mathematics and Computers in Simulation, 21(3): 307–308, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900958.

VanRemoortere:1979:BRCf

[493] P. Van Remoortere. Book review:

Computer networks and simulation: S. Schoemaker (Editor): 1978, 256 pages, Published by North-Holland Publishing Company, Amsterdam, New York, Price: US $35.50/Dfl. 80.00. ISBN 0-444-85208-5. Mathematics and Computers in Simulation, 21 (3):308, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990096X.

VanRemoortere:1979:BRSb

[494] P. Van Remoortere. Book review: Simulation of control systems with special emphasis on modelling and redundancy: I. Troch (Editor): Proceedings of the IMACS symposium of September 27– 29, 1978 in Vienna, 312 pages. North-Holland Publishing Company, Amsterdam and New York, US $41.50 or Dfl. 85.00. ISBN 0-444-85199-2. Mathematics and Computers in Simulation, 21(3):308–309, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900971.

VanRemoortere:1979:BRMe

[495] P. Van Remoortere. Book review: Microprocessors: Design and applications.: A. M. Veronis: Reston Publishing Company, Inc., Reston, Virginia 22090 USA. 272 pages, April 1978. $16.95. Mathematics and Computers in Simulation, 21(3):309–310, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900983.

Anonymous:1979:Bb

[496] Anonymous. Bibliography. Mathematics and Computers in Simulation, 21(3):311–328, November 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900995.

Anonymous:1979:PN

[497] Anonymous. Pages 263–328 (November 1979). Mathematics and Computers in Simulation, 21(3):??, November 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mesnard:1979:SPR

[498] Guy Mesnard. Simulation des phénomènes de réflexion et de transmission des électrons. (French) [Simulation of reflection phenomena and electron transmission]. Mathematics and Computers in Simulation, 21(4): 329–334, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900016.

Benyon:1979:IDS

[499] P. R. Benyon. The inversion of dynamic systems. Mathematics and Computers in Simulation, 21(4):335– 339, December 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900028.

Lefevre:1979:ASD

[500] J. Lefevre and A. Magnus. Accuracy of simple difference-differential equations for blood flow in arteries. Mathematics and Computers in Simulation, 21(4):340–351, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990003X.

Saaty:1979:MMD

[501] Thomas L. Saaty. Mathematical modeling of dynamic decisions; priorities and hierarchies with time dependence. Mathematics and Computers in Simulation, 21(4):352–358, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900041. See

erratum [544].

Liu:1979:NML

[502] Mei-Kao Liu, Thomas C. Myers, and James L. McElroy. Numerical modeling of land and sea breeze circulation along a complex coastline. Mathematics and Computers in Simulation, 21(4):359–367, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900053.

Kincaid:1979:UIM

[503] David R. Kincaid, Roger G. Grimes, and David M. Young. The use of iterative methods for solving large sparse PDE-related linear systems. Mathematics and Computers in Simulation, 21(4):368–375, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900065.

Sigal:1979:UCM

[504] C. E. Sigal, A. A. B. Pritsker, and J. J. Solberg. The use of cutsets in Monte Carlo analysis of stochastic networks. Mathematics and Computers in Simulation, 21(4): 376–384, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900077.

VanRemoortere:1979:BRCg

[505] P. Van Remoortere. Book review: Content addressable parallel processors.: C. C. Foster: Van Nostrand Reinhold, New York, 1976, 233 pages, $11.95. Mathematics and Computers in Simulation, 21(4):385, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900119.

VanRemoortere:1979:BRDd

[506] P. Van Remoortere. Book review: Dynamic programming and stochastic control: D. P. Bertsekas: 1976, 397 pages, Academic Press, New York. Mathematics and Computers in Simulation, 21(4):385, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900090.

VanRemoortere:1979:BRDe

[507] P. Van Remoortere. Book review: Dynamical systems: L. Cesari: 1976, 336 + 314 pages; Academic Press, New York. Mathematics and Computers in Simulation, 21(4):385, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900107.

VanRemoortere:1979:BRDf

[508] P. Van Remoortere. Book review: Dynamic stochastic models from empirical data: R. Kashyap and A. Rao: 1976, 344 pages, Academic Press, New York. Mathematics and Computers in Simulation, 21(4):385, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900120.

VanRemoortere:1979:BRLa

[509] P. Van Remoortere. Book review:

Linear and combinatorial programming: K. G. Murty: 1976, J. Wiley, New York, 310 pp. Mathematics and Computers in Simulation, 21 (4):385, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900144.

VanRemoortere:1979:BRMf

[510] P. Van Remoortere. Book review: Microprogrammable computer architectures: A. B. Salisbury: Published by North-Holland, Amsterdam and N.Y. 1976, 162 pages, US $18.95, ISBN 0-444-00175-1. Mathematics and Computers in Simulation, 21 (4):385, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900156.

VanRemoortere:1979:BRSc

[511] P. Van Remoortere. Book review: System Identification. Advances and Case studies: R. Mehra and D. G. Lainiotis (Editors): Published by Academic Press N.Y. in 1976, 604 pages, US $19.00. ISBN 0-12-487950-0. Mathematics and Computers in Simulation, 21(4):385, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900132.

VanRemoortere:1979:BRSd

[512] P. Van Remoortere. Book review: Simulations-technik, Entwurf und Simulation von Systemen auf digitalen Rechenautomaten: K. Bauknecht, J. Kohlaas and C. A. Zehnder: 1976, 218 pages, Springer-Verlag, Berlin. Mathematics and Computers in Simulation, 21 (4):385, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900089.

VanRemoortere:1979:BRCh

[513] P. Van Remoortere. Book review:

Computer-based medical consultations: MYCIN: E. H. Shortliffe: Published by North-Holland, Amsterdam and N.Y., 1976, 264 pages, US $19.95, ISBN 0-444-00179-4. Mathematics and Computers in Simulation, 21(4): 385–386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900168.

VanRemoortere:1979:BRCi

[514] P. Van Remoortere. Book review:

Computer performance measurement and evaluation methods: analysis and applications: L. Svobodova: Published by North Holland, Amsterdam and N.Y., 1976, 146 pages, US $19.95. ISBN 0-444-00192-. Mathematics and Computers in Simulation, 21(4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990017X.

VanRemoortere:1979:BRDg

[515] P. Van Remoortere. Book review: Digital signal processing: theory design and implementation: A. Peled and B. Liu: 1976, 304 pages, J. Wiley, New York. Mathematics and Computers in Simulation, 21(4):386, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900181.

VanRemoortere:1979:IDT

[516] P. Van Remoortere. Introduction to discrete-time signal processing: S. A. Tretter: 1976, 460 pages, J. Wiley, New York. Mathematics and Computers in Simulation, 21 (4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900193.

VanRemoortere:1979:BRLb

[517] P. Van Remoortere. Book review: Linear estimation and stochastic control: M. H. A. Davis: 1977, 244 pages, Chapman and Hall, London. Mathematics and Computers in Simulation, 21(4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900235.

VanRemoortere:1979:BRLc

[518] P. Van Remoortere. Book review: Linear programming and network flows: M. S. Bazaraa and J. J. Jarvis: 1977, 565 pages J. Wiley, New York. Mathematics and Computers in Simulation, 21(4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900211.

VanRemoortere:1979:BRPc

[519] P. Van Remoortere. Book review: Problem Analysis in Science and Engineering: F. H. Branin and H. Huseyin (Editors): Published by Academic Press N.Y. in December 1977, 528 pages, US $21.50. ISBN 0-12-125550-6. Mathematics and Computers in Simulation, 21(4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900223.

VanRemoortere:1979:PSL

[520] P. Van Remoortere. Problems and solutions in logic design: D. Zissos: 1976, 146 pages, Oxford University Press, London. Mathematics and Computers in Simulation, 21 (4):386, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990020X.

VanRemoortere:1979:BRAc

[521] P. Van Remoortere. Book review: The art and theory of dynamic programming: S. E. Dreyfus, A. M. Law: Published by Academic Press N.Y. in 1977, 310 pages, US $18.50, ISBN 0-12-221860-4. Mathematics and Computers in Simulation, 21(4):387, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900247.

VanRemoortere:1979:BREb

[522] P. Van Remoortere. Book review: Elements of software science: M. H. Halstead: Published by North Holland Amsterdam and N.Y., 1977, 128 pages, US $18.95. ISBN 0-444-00205-7. Mathematics and Computers in Simulation, 21(4):387, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900272.

VanRemoortere:1979:BRSe

[523] P. Van Remoortere. Book review: Software Engineering: K. Gewald, G. Haake and W. Pfadler: Published by R. Oldenburg, München, 1977, 318 pages, 70 figures, 20 tables, DM. 48,00 ISBN 3-486-21491-8. Mathematics and Computers in Simulation, 21(4):387, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900260.

VanRemoortere:1979:BRPd

[524] P. Van Remoortere. Book review: The process view of simulation: W. R. Franta: Published by North Holland, Amsterdam and N.Y., in 1977, 128 pages, US $19.95. ISBN 0-444-00221-9. Mathematics and Computers in Simulation, 21(4):387, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900259.

VanRemoortere:1979:BRF

[525] P. Van Remoortere. Book review: Flow analysis of computer programs: M. S. Hecht: Published by North Holland Amsterdam and N.Y., 1977, 232 pages, US $19.95. ISBN 0-444-00210-3. Mathematics and Computers in Simulation, 21(4):387–388, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900284.

VanRemoortere:1979:BRGa

[526] P. Van Remoortere. Book review:

“GPSS-Fortran”, Einführung in die Simulation diskreter Systems mit Hilfe eines Fortran-Programmpaketes: B. Schmidt: 1977, 298 pages, Springer-Verlag, Berlin. Mathematics and Computers in Simulation, 21(4):388, December 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900338.

VanRemoortere:1979:BRMg

[527] P. Van Remoortere. Book review: Micro-computer handbook: C. J. Sippl: 1977, 454 pages, Petrocelli/Charter, New York. Mathematics and Computers in Simulation, 21 (4):388, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847547990034X.

VanRemoortere:1979:BRMh

[528] P. Van Remoortere. Book review:

Mathematical modeling and digital simulation for engineers and scientists: J. M. Smith: 1977, 332 pages, J. Wiley, New York. Mathematics and Computers in Simulation, 21 (4):388, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900351.

VanRemoortere:1979:BRMi

[529] P. Van Remoortere. Book review: Mathematical Models for the study of the reliability of systems: A. Kaufmann, D. Groucho and R. Groun: Published by Academic Press N.Y. in 1977. 236 pages, US $18.50, ISBN 0-12-402370-3. Mathematics and Computers in Simulation, 21(4):388, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900302.

VanRemoortere:1979:BRMj

[530] P. Van Remoortere. Book review: Méthodes et techniques de traitement du signal et applications aux mesures physiques: J. Max: 1977, 379 pages, Masson, Paris. Mathematics and Computers in Simulation, 21(4):388, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900314.

VanRemoortere:1979:BRQ

[531] P. Van Remoortere. Book review: Qualitative analysis of large scale dynamical systems: Michel A. N. and Miller R. K.: 1977, 289 pages, Academic Press, New York. Mathematics and Computers in Simulation, 21 (4):388, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900326.

VanRemoortere:1979:BRU

[532] P. Van Remoortere. Book review: Using the computer in the social sciences. A non-technical approach: R. J. Hy: Published by North Holland, Amsterdam and N.Y., 1977, 156 pages, US $12.95. ISBN 0-444-00219-7. Mathematics and Computers in Simulation, 21(4):388, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900296.

VanRemoortere:1979:BRAd

[533] P. Van Remoortere. Book review:

AGARD Conference Proceedings No. 249 on Piloted Aircraft Environment Simulation Techniques: AGARD-CP 249: 1978. AGARD, NATO, 7, rue Ancelle, F-92200 Neuilly sur Seine, France. ISBN 92-835-0222-1. 26 papers. Mathematics and Computers in Simulation, 21(4):388–389, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900363.

VanRemoortere:1979:BRGb

[534] P. Van Remoortere. Book review: Gaussian random processes (Translated from Russian by A. B. Aries): I. A. Ibragimov and Y. A. Rozanov: Published by Springer-Verlag, Berlin in 1978, 280 pages, DM 49.60, ISBN 3-540-90302-X. Mathematics and Computers in Simulation, 21(4):389, December 1979. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900399.

VanRemoortere:1979:BRLd

[535] P. Van Remoortere. Book review:

Locally interacting systems and their application in Biology. Proceedings of the School-Seminar on Markov interaction processes in biology, held in Pushchino, Moscow in March 1976: R. L. Dobrushin, V. I. Kryukov, A. L. Toom(Editors). Published by Springer-Verlag, Berlin in 1978, 202 pages, DM 24.80, ISBN 3-540-08450-9. Mathematics and Computers in Simulation, 21(4):389, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900375.

VanRemoortere:1979:BRN

[536] P. Van Remoortere. Book review: Numerical analysis for computer science: I. A. Dodes: Published by North Holland Publishing Company, Amsterdam and N.Y., in 1978, 640 pages, US $19.95 ISBN 0-0-444-00238-3. Mathematics and Computers in Simulation, 21(4):389, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900387.

VanRemoortere:1979:BRIb

[537] P. Van Remoortere. Book review: Introduction of Simulation and SLAM: A. A. B. Pritsker and C. D. Pegden: Distributed by Halsted Press, a division of John Wiley and Sons, Inc. N.Y. and Systems Publishing Corporation, P.O. Box 2161, West Lafayette, Indiana 47906 USA. 1979, 588 pages, US$25, ISBN 0-470-26588-4. Mathematics and Computers in Simulation, 21(4):390, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900429.

VanRemoortere:1979:BRR

[538] P. Van Remoortere. Book review:

Realzeitrechner-Modelle: B. Walke: Published by R. Oldenburg Verlag, München, 1978, 240 seiten, 81 Abbildungen, 8 Tabellen, 5 ALGOL-Programme, DM.44,00. ISBN 3-486-21551-5. Mathematics and Computers in Simulation, 21(4):390, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900417.

VanRemoortere:1979:BRIc

[539] P. Van Remoortere. Book review: An introduction to the general theory of algorithms: M. Machtey and P. Young: Published by North Holland, Amsterdam and N.Y., 1978, 272 pages, US $19.95. ISBN 0-444-00226-X. Mathematics and Computers in Simulation, 21(4):390, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900405.

Anonymous:1979:BEP

[540] Anonymous. Bibliography: Edited by P. Van Remoortere. Mathematics and Computers in Simulation, 21 (4):391–407, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475479900430.

Anonymous:1979:AIV

[541] Anonymous. Author index: Volume XXI (1979). Mathematics and Computers in Simulation, 21(4):408– 409, December 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900442.

Anonymous:1979:SIV

[542] Anonymous. Subject index: Volume XXI (1979). Mathematics and Computers in Simulation, 21 (4):410, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475479900454.

Anonymous:1979:PD

[543] Anonymous. Pages 329–410 (December 1979). Mathematics and Computers in Simulation, 21(4):??, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1981:EMM

[544] Anonymous. Erratum: Mathematical modeling of dynamic decisions; priorities and hierarchies with time dependence: T. L. Saaty, Mathematics and Computers in Simulation XXI (1979) 352– 358. Mathematics and Computers in Simulation, 23(1):98, March 1981. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900197. See

[501].